### v1

# LoSiA: Efficient High-Rank Fine-Tuning via Subnet Localization and Optimization

Abstract
Parameter-Efficient Fine-Tuning (PEFT) methods, such as LoRA, significantly reduce the number of trainable parameters by introducing low-rank decomposition matrices. However, existing methods perform extensive matrix multiplications in domain specialization tasks, resulting in computational inefficiency and sub-optimal fine-tuning performance. Hence, we propose LoSiA [^1] The source code is released at https://github.com/KlozeWang/LoSiA. (Low-Resources Subnet Integration Adaptation), an innovative method that dynamically localizes and optimizes critical parameters during the training process. Specifically, it identifies a sub-network using gradient sparsity analysis and optimizes it as the trainable target. This design enables effective high-rank adaptation by updating only the sub-network parameters, reducing the additional matrix multiplication. We also present LoSiA-Pro, a faster implementation of LoSiA, which reduces the training latency by about $27\%$ compared to LoRA. Extensive evaluations show that our method achieves minimal performance drop compared to full fine-tuning, while requiring the least training time across domain specialization and common-sense reasoning tasks. Further analysis shows that LoSiA also reduces forgetting during continued training.
LoSiA: Efficient High-Rank Fine-Tuning via
Subnet Localization and Optimization
Xujia Wang  Yunjia Qi  Bin Xuthanks: Corresponding author. Tsinghua University Department of Computer Science and Technology wang-xj22@mails.tsinghua.edu.cn
1 Introduction
Figure 1: Overview of LoSiA. It locates and optimizes core sub-network in asynchronous periods.
Large language models, when fine-tuned via supervised learning, can be effectively adapted to downstream tasks such as mathematics Shao et al. (2024), programming Hui et al. (2024), and domain knowledge reasoning Wei et al. (2021). Although full parameter fine-tuning often yields the best performance, updating billions of parameters is computationally expensive and resource-intensive. To address this, parameter-efficient fine-tuning (PEFT) updates only a small subset of parameters to reduce GPU memory usage and communication overhead while maintaining performance comparable to full fine-tuning Houlsby et al. (2019); Ding et al. (2023).
Among PEFT approaches, LoRA Hu et al. (2022) has gained widespread adoption by introducing low-rank matrices to approximate full weight updates, enabling competitive performance with significantly reduced computational and economic costs Taori et al. (2023). Variants in the LoRA family further refine the method by biased fine-tuning modules Zhu et al. (2024); Hayou et al. (2024a) or dimensions Meng et al. (2024a) to accelerate convergence and achieve superior performance. However, constrained by the low-rank assumption, these paradigms often struggle to balance model performance and efficiency, particularly in domain-specific tasks Yang et al. (2024); Ghosh et al. (2024) and continual learning scenarios Shuttleworth et al. (2024a). In such settings, low rank configurations (e.g., $8$ or $16$) can lead to performance degradation and underfitting Biderman et al. (2024). Although increasing the rank may mitigate these issues, it introduces additional memory consumption, extensive floating point operations, and risks of overfitting or convergence difficulties Kalajdzievski (2023); Borse et al. (2024). Recent studies have attempted to approximate high-rank updates by accumulating multiple low-rank components. However, these approaches still suffer from issues such as locally low-rank updates Meng et al. (2024b); Lialin et al. (2023) or increased computational complexity Zhao et al. (2024a). Therefore, while the low-rank assumption offers notable improvements in efficiency, it also introduces inherent limitations.
The Lottery Ticket Hypothesis Frankle and Carbin (2019) suggests that dense neural networks contain trainable sub-networks capable of achieving comparable test accuracy. This prompts us to reconsider the route of PEFT and explore an alternative: Can we identify and fine-tune such sub-networks within the backbone model to achieve high-quality adaptation more efficiently?
To answer this question, we propose LoSiA (Low-Resources Subnet Integration Adaptation), a novel PEFT framework that dynamically localizes and optimizes critical sub-networks periodically, as illustrated in Figure 1. LoSiA asynchronously selects a core sub-network for each layer by calculating sensitivity-based importance scores and performing greedy selecting algorithms. Following localization, it fine-tunes the identified sub-network and applies a rewarming learning rate strategy to promote stable and robust training. The design enables real-time high-rank updates without introducing additional matrix multiplication overhead, which reduces training latency and ensures no steep increase in training time for high-rank updating. Additionally, LoSiA does not introduce extra architectural components and only requires an optimizer replacement for seamless deployment. Extensive experiments demonstrate its superior performance among PEFT baselines on domain-specific, commonsense reasoning tasks, while mitigating forgetting in continue learning. We also propose LoSiA-Pro, a more refined equivalent implementation of LoSiA, which significantly reduces the activation storage and computational complexity in backward propagation. LoSiA-Pro speeds up training 1.38×1.38$\times$1.38 × compared to LoRA and 2.68×2.68$\times$2.68 × compared to DoRA.
In summary, our contributions are as follows.
(1) Innovatively introduces the structure of the sub-network to the field of parameter-efficient fine-tuning. We develop periodic subnet localization, optimization and integration techniques to dynamically capture and adapt task-essential parameters.
(2) We propose LoSiA, a novel PEFT approach that dynamically localizes and optimizes sub-networks to achieve high performance, efficiency, and usability. To further improve practicality, we also introduce a faster variant, LoSiA-Pro.
(3) We conduct extensive evaluations across multiple models and benchmarks. LoSiA outperforms all advanced PEFT baselines on domain-specific and commonsense reasoning tasks, while also accelerating training by 1.15×1.15$\times$1.15 × compared to LoRA. Moreover, its efficient variant, LoSiA-Pro, achieves a further speedup of 1.38×1.38$\times$1.38 ×.
2 Related Work
Parameter-Efficient Fine-Tuning
Full parameter fine-tuning (FFT) adapts pre-trained models to downstream tasks by updating all model parameters Wei et al. (2022), but it is often computationally expensive. In contrast, parameter-efficient fine-tuning (PEFT) methods update only a small subset of parameters, significantly reducing training costs while still maintaining strong performance. LoRA Hu et al. (2022) approximates parameter updates as the product of low-rank matrices, achieving promising performance for tasks such as instruction tuning Ghosh et al. (2024). Enhanced variants such as PiSSA Meng et al. (2024a) accelerate convergence by prioritizing major singular vectors, while DoRA Liu et al. (2024) improves performance in low-rank by decomposing updates into directional and magnitude components. Other derivatives such as LoRA+ Hayou et al. (2024b), LoRA-GA Wang et al. (2024a) and LoRA-Dash Si et al. (2025) refine the framework by directional or module biased fine-tuning.
However, recent studies Jiang et al. (2024b); Biderman et al. (2024); Ghosh et al. (2024) reveal that the low-rank structure limits effectiveness in knowledge-intensive domains (e.g., mathematics, coding). Advanced solutions adopt novel strategies: 1) Architectural modifications through MoE-based LoRA combinations Zadouri et al. (2023); Li et al. (2024); Wang et al. (2024b) for multitask scenarios; 2) High-rank fine-tuning via accumulated low-rank projections, such as ReLoRA Lialin et al. (2023), MoRA Jiang et al. (2024a) and GaLore Zhao et al. (2024a) to enhance training effectiveness. However, these approaches incur an increase in architectural complexity or a drop in throughput. Rare methods achieve an optimal balance between performance, training latency, and implementation simplicity.
Skill Localization and Pruning
LLM pruning reduces neural network size by eliminating redundant or less critical parameters. Previous work demonstrates that sparse networks can play crucial roles Frankle and Carbin (2019); Yao et al. (2025). Panigrahi et al. (2023) identifies critical parameters in fine-tuned LMs by optimizing masks of grafted models, although such methods require additional training time and data. Alternatively, gradient- and sensitivity-based metrics enable real-time identification of task-aware parameters Molchanov et al. (2019); Sanh et al. (2020); Zhang et al. (2022). Recent advances adapt the techniques to PEFT: Zhang et al. (2023) prunes LoRA trainable parameters, while Feng et al. (2024) applies the approach to continual learning scenarios.
3 Method
Definition
[…]
3.1 Structure of Gradients
Inspired by pruning techniques, we minimize the mean squared error (MSE) $\mathcal{L}_{\text{MSE}}$ between the outputs of models $f_{k},f^{\prime}_{k}$, where both are trained on $f_{k-1}$ with trainable parameter set $P_{0}$ and $P$, respectively. For SGD optimizers, we derive the bound:
$\scriptstyle\mathcal{L}_{MSE}$ $\leq\eta^{2}\frac{\|1_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B}_{k})\|^{2}_{F}\|x\|^{2}_{F}}{M}$ (1)
For AdamW optimizers, $\mathcal{L}_{MSE}$ admits a similar bound through $\nabla W$ in most cases (Appendix A.1.1). Thus, gradient magnitudes in $P$ provide an upper bound for the approximation error, while prioritizing larger gradients yields tighter bounds. We therefore seek optimal subsets $P$ to capture the parameters with large gradient magnitudes.
Ideally, selecting the $top\text{-}K$ entries of $\nabla W$ would suffice, but storing and fine-tuning sparse matrices compromises efficiency. Instead, we claim that a suitable pattern for $P$ corresponds to the parameters in subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$, i.e., all connections between the input neuron set $X_{S}$ and the output neuron set $Y_{S}$.
Figure 2: Gradient Magnitude Distribution of proj_v. Large gradients follow a sparse subnet distribution.
To validate this selection paradigm, Figure 2 visualizes the gradient magnitude distributions in LLaMA-2 7B’s proj_v layer: The 32 attention heads exhibit a significant disparity in attention scores, while gradients strongly correlate with output neurons $Y_{S}$. In particular, a consistent subset of input neurons $X_{S}$ (green markers, x-axis) contributes dominantly to all attention heads. The sparse pattern remains consistent in MLP layers (Appendix A.2.1). We therefore focus on fine-tuning subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$ termed the core subnet - rather than the entire network.
3.2 Subnet Localization
To localize core subnets efficiently, an ideal algorithm should satisfy three key requirements: 1) Efficiency: no additional data requirements or significant latency. 2) Lightweight: minimal GPU memory overhead. 3) Dynamic Awareness: enable real-time localizations during training. To meet these objectives, the subnet localization process is divided into two stages:
Parameter Importance Calculation
To quantify parameter importance $I(\cdot)$, existing approaches LeCun et al. (1989); Ma et al. (2023) observe the change in loss assuming $W_{k}=0$ for the $k$-th parameter. Adopting second-order Taylor expansion, element-wise $I(\cdot)$ is estimated as:
[…]
Here, $H$ stands for the Hessian matrix. However, Eq.2 is difficult to calculate in real-time. We derive a micro-batch approximation:
[…]
Furthermore, estimates by single micro-batch may introduce bias by ignoring training dynamics. Sensitivity smoothing and uncertainty quantification Zhang et al. (2022) are used to handle the problem. For the training step $i$, compute an exponential moving average (EMA) $\overline{I}_{i}$ for $I_{i}$ ,and uncertainty $\overline{U}_{i}$ for variation $\Delta I_{i}=|I_{i}-\overline{I}_{i}|$:
$\overline{I}_{i}(W_{k})$ $=\beta_{1}\overline{I}_{i-1}(W_{k})+(1-\beta_{1})I_{i}(W_{k})$ (4)
$\overline{U}_{i}(W_{k})$ $=\beta_{2}\overline{U}_{i-1}(W_{k})+(1-\beta_{2})|\Delta I_{i}(W_{k})|$ (5)
$s(W_{k})$ $=\overline{I}(W_{k})\cdot\overline{U}(W_{k})$ (6)
where $\beta_{1},\beta_{2}\in(0,1)$ is the EMA factors. We regard $s(\cdot)$ as a appropriate importance assessment. Notably, to obtain gradients $W$ asynchronously, LoSiA use per-layer weight updates Lv et al. (2024), executing the optimization during backpropagation without storing gradients.
Core Subnet Localization via Importance Scores
For a subnet $S$ selected from origin network $S_{0}=(\{i\}_{i=1}^{n},\{j\}_{j=1}^{m},W)$, define its importance as:
$s(S)=\sum_{i\in X_{S}}\sum_{j\in{Y_{S}}}s(W_{ij})$ (7)
The objective is to identify optimal subset $S$ and maximize $s(S)$, while satisfying the memory constraint $\max\{\frac{|X_{S}|}{n},\frac{|Y_{S}|}{m}\}\leq p$, where $p\in(0,1)$ represents the rank factor. However, the task is NP-Hard. Leveraging observations in Section 3.1 regarding gradient magnitude sparsity patterns, we develop greedy selection algorithms to identify critical input and output neuron sets $(X_{S},Y_{S})$:
Algorithm 1 Greedy Strategy for Localization
1:function Row2Column($q$)
2:     $\text{sums}\leftarrow\text{Sum}(q,\text{dim}=1)$
[…]
4:     $\text{sums}\leftarrow\text{Sum}(q[\text{rows},:],\text{dim}=0)$
[…]
7:end function
Algorithm 1 implements a row-major greedy approach that selects the $top\text{-}K$ rows by importance score summation, and then finds $\{Y_{S}\}$ to maximum $s(S)$ given fixed $\{X_{S}\}$. A column-major greedy algorithm Column2Row is also considered. The better selection result of two strategies is selected.
Dimensionality Reduction in Output Layer Fine-Tuning
While prior work Chen et al. (2024) has established the benefits of fine-tuning the output layer in conjunction with PEFT methods, the approach remains computationally prohibitive for large-vocabulary models (e.g., Gemma-2B). However, empirically, backward propagation through the output layer exhibits gradient sparsity, with only a limited subset of tokens receiving significant updates. Building on this insight, LoSiA easily implements an efficient optimization strategy by constructing a tunable subnet $S=(X_{S_{0}},Y_{S},W_{X_{S_{0}},Y_{S}})$ of the output layer, where $|Y_{S}|=p_{o}|Y_{S_{0}}|$, and $p_{o}\in(0,1)$ denoting the dimension reduction factor.
3.3 Subnet Optimization and Intergration
During fine-tuning, the locations of core subnets may undergo dynamic shifts, as illustrated in Figure 3. Although a small subset of neurons is consistently selected, peripheral components exhibit significant temporal variability. Fine-tuning with a static subnet risks model underfitting and neuron over-specialization. To address the issue, we introduce an asynchronous and periodic subnet re-localization mechanism that adapts to the evolving network topology.
Figure 3: Core Subnet Distribution during Training. The subnet various across different training iterations.
Naive periodic re-localization strategies can induce training instability and loss spikes Lialin et al. (2023). Furthermore, the storage requirements for $\overline{I}(\cdot),\overline{U}(\cdot)$ in a synchronous period would lead to a scaling of GPU memory overhead. Therefore, we propose asynchronous periodic localization coupled with rewarmings of learning rate. Consider a model $f$ with $L$ decoder layers $\{D_{l}\}_{l=0}^{L-1}$, where each layer $D_{l}$ contains $K$ linear layers $\{W_{l,k}\}_{k=1}^{K}$ with corresponding core subnets $\{S_{l,k}\}_{k=1}^{K}$. The training timeline is divided into time slots of $T$ steps, such that for time slots $[iT,(i+1)T),i=1,2,\ldots$, we:
  1. 1.
    Compute $\overline{I}(\cdot),\overline{U}(\cdot)$ for layer $D_{l}$ in the time slot $[(kL+l-1)T,(kL+l)T),k\in\mathbb{N}$.
  2. 2.
    Sequentially reselect $S_{l}$ by $s(\cdot)$ before step $t=(kL+l)T$, the end of time slots in 1.
This yields a reselection period of $\overline{T}=LT$ for each layer. The asynchronous design ensures that at any step, there is exactly one decoder layer calculating $\overline{I}(\cdot),\overline{U}(\cdot)$ and one rewarming, which greatly reduces additional GPU memory overheads for importance score calculation. We also implement a learning rate rewarm-up to further enhance training stability. Formally, the learning rate at step $t$ is:
$\overline{lr}(t)=\begin{cases}\frac{t-(kL+l)T}{T}\cdot\text{lr}(t)&\text{if C}\\ \text{lr}(t)&\text{otherwise}\\ \end{cases}\$ (8)
The condition $C$ is $t\in[(kL+l)T,(kL+l+1)T)\text{ and }t>T_{w}$, where $T_{w}$ is the warmup duration. Figure 4 illustrates the timelines of importance calculation and the learning rate rewarming across layers, with re-localization sandwiched between them.
Figure 4: Asynchronous Periodic Subnet Reselection and Learning Rate Rewarming Mechanism (in a 5-layer model for example).
3.3.1 Faster Implementation (LoSiA-Pro)
Through subnet fine-tuning, LoSiA can further mitigate activation storage and backward latency. The gradient for the subnet $S$ can be factorized as:
$\begin{split}\frac{\partial\mathcal{L}}{\partial W_{S}}&=\frac{\partial\mathcal{L}}{\partial W}[X_{S},:][:,Y_{S}]\\ &=(x^{T}[X_{S},:])(\frac{\partial\mathcal{L}}{\partial y}[:,Y_{S}])=\tilde{L}_{S}\tilde{R}_{S}\end{split}$ (9)
Noticing $\tilde{L}_{S}\in\mathbb{R^{np\times bs}},\tilde{R}_{S}\in\mathbb{R}^{bs\times mp}$, the input activation storage is reduced by a factor $p$, while the computational complexity of gradient calculation is reduced from $O(nmbs)$ to $O(nmbsp^{2})$. We named the method LoSiA-Pro, a refined equivalent implementation of LoSiA. It offers a 27.6% latency reduction compared to LoRA with Gradient Check-Pointing, while reducing 13.4GB GPU memory consumption compared to LoSiA training without Gradient Check-Pointing.
4 Experiments
We evaluate LoSiA across a broad range of model scales and datasets, conducting rigorous comparisons with common baselines. On both domain-specific and common-sense reasoning tasks, the method demonstrates robust performance with significantly reduced training overheads. The experiments highlight that LoSiA effectively promotes both training efficiency and task proficiency.
4.1 Experimental Setup
Datasets
Models are trained on downstream tasks in the domains of mathematics, coding, and general capabilities. Specifically, training sets are sampled by 50,000 random entries from MetaMathQA, Magicoder, and Alpaca-GPT4, respectively. The GSM8K, MBPP, and MMLU benchmark are for testing. Additionally, we also compared LoSiA with baseline methods on eight common sense reasoning tasks. More details regarding the datasets can be found in the Appendix.
Table 1: Comparison of PEFT Methods Across Models on Domain-Specific Tasks. Accuracy is reported, alongside with memory consumption (GB), maximum per-task training time (h). The numbers in parentheses indicate the training time of LoSiA-Pro, which is a refined and computationally equivalent implementation of LoSiA.
Model Method Mem(GB) Time(h) GSM8K MBPP MMLU Avg.
5-shot 0-shot,CoT Pass@1 Pass@10 0-shot,PPL 5-shot,GEN
Gemma 2B FFT $50.1$ $11.0$ $46.4$ $50.4$ $33.0$ $43.4$ $36.1$ $37.0$ $41.05$
LoRA $36.1$ $15.0$ $35.7$ $41.1$ $26.0$ $36.6$ $34.9$ $31.2$ $34.25$
PiSSA $36.1$ $14.9$ $38.5$ $46.5$ $26.4$ $39.0$ $33.8$ $32.6$ $36.13$
DoRA $37.3$ $29.8$ $39.7$ $43.0$ $31.4$ $\mathbf{43.2}$ $36.2$ $37.1$ $38.43$
GaLore $37.5$ $14.2$ $39.3$ $44.7$ $\mathbf{31.6}$ $42.6$ $36.6$ $35.5$ $38.38$
LoSiA (-Pro) $36.9$ $\mathbf{10.3\ (9.4)}$ $\mathbf{42.8}$ $\mathbf{49.7}$ $30.7$ $43.0$ $\mathbf{37.5}$ $\mathbf{37.4}$ $\mathbf{40.18}$
LLaMA 2-7B FFT $64.1$ $27.5$ $46.6$ $46.9$ $29.9$ $40.2$ $45.2$ $42.5$ $41.88$
LoRA $23.7$ $33.3$ $42.9$ $46.7$ $26.0$ $37.8$ $42.3$ $37.3$ $38.83$
PiSSA $23.7$ $33.1$ $43.5$ $46.2$ $26.8$ $36.6$ $42.7$ $38.5$ $39.05$
DoRA $24.2$ $68.3$ $\mathbf{45.0}$ $\mathbf{47.2}$ $26.0$ $34.4$ $44.1$ $36.7$ $38.90$
GaLore $23.7$ $39.6$ $42.2$ $45.3$ $28.0$ $39.0$ $43.1$ $41.2$ $39.80$
LoSiA (-Pro) $21.9$ $\mathbf{26.1\ (21.8)}$ $44.7$ $46.7$ $\mathbf{28.4}$ $\mathbf{39.4}$ $\mathbf{45.0}$ $\mathbf{41.5}$ $\mathbf{40.95}$
LLaMA 2-13B FFT-8Bit $77.1$ $53.0$ $61.2$ $55.7$ $35.7$ $43.2$ $53.6$ $56.2$ $50.93$
LoRA $36.9$ $56.0$ $58.6$ $\mathbf{56.4}$ $34.1$ $44.8$ $52.6$ $53.7$ $50.03$
PiSSA $36.9$ $55.5$ $53.4$ $55.2$ $34.5$ $44.8$ $52.0$ $48.8$ $48.11$
LoSiA (-Pro) $36.9$ $\mathbf{46.5\ (38.6)}$ $\mathbf{59.0}$ $54.0$ $\mathbf{34.9}$ $\mathbf{48.2}$ $\mathbf{53.1}$ $\mathbf{55.7}$ $\mathbf{50.82}$
Implementation Details
We employ Gemma 2B, LLaMA-2 7B, and LLaMA-2 13B as the backbone models. The effectiveness of LoSiA is evaluated against parameter-efficient fine-tuning (PEFT) baselines, namely LoRA, DoRA, PiSSA, and GaLore. For control of consistency in memory consumption, the rank $r$ of LoRA, DoRA, and PiSSA is set to 64. For GaLore, the gradient projection rank $R$ is set to 512 with the full projection strategy. In the case of LoSiA, the rank factor $p$ is set to $\frac{1}{8}$. The learning rate is $6\times 10^{-5}$ for MetaMathQA and $5\times 10^{-5}$ for the rest, with time slots $T$ of $100$ for MetaMathQA and $150$ for the rest.
Additionally, both GaLore and LoSiA incorporate the output layer into the fine-tuning process. Dimension reduction factor $p_{0}$ is set to $\frac{1}{64}$ for Gemma 2B, $\frac{1}{8}$ for LLaMA-2 7B, and $1$ for LLaMA-2 13B in LoSiA. The PEFT modules are applied to all linear layers within the transformer. The training batch size is set to 4, the warm-up ratio is set to 0.1 and the model is trained by 3 epochs. For training stability, the backbone model are in precision of BF16 and low-rank modules are upcasted to FP32.  [^2] Trained with LLaMA-Factory Zheng et al. (2024). Upcasting to FP32 only costs an additional 0.6GB of memory and trains faster than BF16 in practice. All of the experiments are conducted on single NVIDIA A800 80GB GPU. Further details (including implementation details for common-sense reasoning tasks) can be found in Appendix A.3.
Table 2: Comparison of PEFT Methods on Commen-Sense Reasoning Tasks, using LLaMA-2 7B as the backbone model. Evaluations are PPL-based in lm-evaluation-harness and we report the ACC metric.
Method Mem(GB) Time(h) ARC-C ARC-E HellaSwag Winogrande PIQA OBQA SIQA BoolQ Avg.
LoRA $19.46$ $10.0$ $50.28$ $79.71$ $59.86$ $73.88$ $79.33$ $55.00$ $56.86$ $\mathbf{88.07}$ $67.87$
PiSSA $19.18$ $10.4$ $51.19$ $79.80$ $62.36$ $77.74$ $80.41$ $56.60$ $59.88$ $87.71$ $69.46$
DoRA $20.42$ $25.6$ $51.71$ $79.34$ $59.86$ $\mathbf{79.24}$ $79.98$ $59.60$ $59.57$ $88.04$ $69.67$
GaLore $\mathbf{18.24}$ $16.7$ $48.63$ $79.97$ $60.07$ $76.24$ $80.09$ $56.80$ $56.65$ $82.60$ $67.63$
LoSiA $18.68$ $\mathbf{9.2}$ $\mathbf{52.22}$ $\mathbf{80.26}$ $\mathbf{65.05}$ $77.19$ $\mathbf{81.50}$ $\mathbf{61.40}$ $\mathbf{61.05}$ $84.13$ $\mathbf{70.35}$
Figure 5: Overheads Comparison of PEFT methods training with and without Gradient Check-Pointing (GC). Taking training arguments in Table 1 as example.
4.2 Main Results
Table 1 presents the performance of LoSiA compared to baseline methods across Gemma-2B, LLaMA2-7B, and LLaMA2-13B models. For GSM8K, we report 0-shot Chain-of-Thought (CoT) and 5-shot accuracy to reveal the model’s reasoning capability and few-shot prompting performance. For MBPP, we report the Pass@1 and Pass@10 metrics. For MMLU, we report both 5-shot generation and perplexity-based results. The metrics are intended to measure the quality of generation and knowledge proficiency, respectively. Table 2 shows results on common-sense reasoning tasks, extracting the option with minimum perplexity and reporting ACC metric following lm-evaluation-harness. The test setup provides a robust measure of intrinsic knowledge acquisition.
LoSiA effectively reserves knowledge
LoSiA demonstrates superior knowledge retention, as evidenced by perplexity-based evaluations. It outperforms LoRA by 2.48% on commonsense reasoning tasks and maintains an average 1.93% improvement on MMLU (0-shot, PPL). Unlike low-rank methods, LoSiA’s sparse, high-rank fine-tuning approach enables localized knowledge retention while shifting likelihood toward correct answers.
LoSiA demonstrates superior performance in generalization
In domain-specific tasks, LoSiA achieves average improvements of 1.75%, 1.15%, and 0.79% compared to the best baseline, respectively. High-rank update methods such as GaLore also exhibit relatively stable performance. The method shows its strength in problem-solving metrics (GSM8K, MBPP Pass@1, and MMLU 5-shot), suggesting that LoSiA provides strong generalization capabilities by applying learned knowledge to address various problems. Notably, while performing comparable to Full-Parameter Fine-Tuning (FFT) with only 0.64% of degradation in average, LoSiA significantly reduces computational resources, highlighting its practical efficiency.
LoSiA and LoSiA-Pro greatly improve training efficiency
Figure 5 compares the training overheads of various PEFT methods. Contrast to baselines such as DoRA which incur significant additional FLOPs, LoSiA shows superior efficiency in both training time and memory usage. By eliminating extra matrix multiplication operations, LoSiA achieves faster training speeds. Its refined implement, LoSiA-Pro, further compresses activation storage by at least 22.8GB (w GC) and reduce the training time up to 34% (w/o GC) compared to LoRA by saving and computing on partial activations. A detailed training latency and GPU memory measurement is in Appendix A.4.
4.3 Ablation Study
This section assesses the functionality of sensitivity importance-aware localization, asynchronous mechanism, and re-warmups, alongside with robustness analysis of LoSiA. We present comprehensive ablation studies in Table 3 and training dynamics in Figure 6. Additional robustness tests for rank factor selection are provided in the Appendix.
Table 3: Ablation Study of LoSiA on GSM8K and MMLU, using LLaMA-2 7B as the backbone.
Model GSM8K MMLU Avg.
Vanilla LoSiA $\mathbf{44.66}$ $\mathbf{44.95}$ $\mathbf{44.81}$
Synchronous Localization (SL) $42.76$ $44.13$ $43.45$
Gradient-based Localization (GL) $43.00$ $44.88$ $43.94$
w/o Warm-up during Selection (WDS) $38.06$ $44.21$ $41.14$
w FFT lm_head (FFTO) $43.96$ $44.32$ $44.14$
Figure 6: Loss Curves of Baselines and LoSiA Variants, training on MetaMathQA and Alpaca-GPT4.
Asynchronous mechanism yields more stable training
Variant SL refers to using a synchronous layer-wise localization mechanism. However, it causes loss fluctuation, destabilizes later training, and degenerate the model performance by 1.36% in average, while asynchronous updates produce more stable loss curves.
Sensitivity-based Importance versus Gradient-based Importance
Variant GL uses absolute gradients as the importance score. On MMLU, its performance remains comparable but is biased toward Humanities tasks (see Table 11), while its accuracy on GSM8K drops by 1.66%. Sensitivity-based scores, which aggregate multi-sample information, are more effective to capture general patterns in linear layers compared to biased gradients. However, gradient-based LoSiA exhibits promising results. In practice, the storage of $\overline{I}(\cdot),\overline{U}(\cdot)$ (about 1GB memory occupation on LLaMA-2 7B) can be eliminated using gradient-based importance if needed. Further discussion is provided in Appendix A.2.2.
Effect of rewarming and full fine-tuning the output layer
The variant w/o WDS, which omits rewarm-ups, introduces instability of the loss, leads to under-fitting and ultimately impairs final performance. w FFTO fully fine-tunes the output layer, shows a performance comparable to LoSiA with additional trainable parameters. It highlights the effectiveness of extracting tunable subnets on the output layer in LoSiA. In permissible GPU memory constraints, fully training the output layer shows promising performance and is also recommended.
Table 4: Robustness of Time Slot $T$ Across a Series of Data Scales. Trained by MetaMathQA and evaluated by GSM8K on LLaMA-2 7B.
Method @30K @50K @70K
LoRA $41.39$ $42.86$ $44.58$
$T$ LoSiA
$25$ $\mathbf{42.99}$ $\mathbf{43.37}$ $42.07$
$50$ $\mathbf{42.91}$ $42.46$ $42.15$
$75$ $41.09$ $\mathbf{44.05}$ $\mathbf{47.46}$
$100$ $40.49$ $\mathbf{44.66}$ $\mathbf{46.17}$
$125$ $39.88$ $42.23$ $\mathbf{45.19}$
$150$ $39.12$ $40.41$ $42.84$
Robustness across varying data scales and time slot lengths
Table 4 assesses the performance of LoSiA across different training data scales. LoSiA consistently outperforms LoRA, demonstrating stability and robustness. Furthermore, the optimal time slot $T$ is positively correlated with the size of training set, while LoSiA shows transcendent performance within a reasonable range of $T$.
4.4 Analysis
Selection Distribution
We analyze the core subnet selection frequency of neurons in Figure 7. The frequently selected neurons remain similar under different rank factor $p$, while smaller $p$ produces more concentrated distribution patterns. This indicates that LoSiA effectively identifies and optimizes critical neurons with limited training budgets, while simultaneously adjusting marginal parameters to further enhance generalization capability.
Reduce Intruder Dimensions
[…]
Figure 7: Selected Frequency Distributions of Neurons in Core Subnets. Sorted frequencies are ploted in black.
Figure 8: Similarities of Top-500 Largest Singular Vectors between Pre- and Post-Fine-Tuning Weights.
To evaluate LoSiA’s efficacy in continual learning, we perform sequential fine-tuning on Hellaswag, PiQA, BoolQ, SiQA, and Winogrande datasets on LLaMA-2 7B. We employs Average Performance (AP) Chaudhry et al. (2018), Forward Transfer (FWT) Lopez-Paz and Ranzato (2017), and Backward Transfer (BWT) Ke and Liu (2023) to assess overall performance, knowledge transfer ability from previous tasks to current task, and level of forgetting, respectively. Details of the metrics and experiments are provided in A.3.4.
Table 5: Results of Continue Learning with Sequential PEFTs on Five Commen-Sense Reasoning Tasks.
Method AP($\uparrow$) FWT($\uparrow$) BWT($\uparrow$)
Seq-LoRA $66.62$ $\mathbf{1.46}$ $-8.04$
Seq-LoSiA $\mathbf{70.48}$ $-0.20$ $\mathbf{-3.54}$
The results in Table 5 demonstrate that LoSiA outperforms LoRA in mitigating forgetting with 4.5% in BWT and achieves a 3.86% improvement in average performance of sequential fine-tuning. This aligns with our hypothesis that LoSiA exhibits stronger robustness in continue learning, indicating that our method can adapt to more diverse application scenarios than existing baselines.
5 Conclusion
We present LoSiA, a novel PEFT framework that dynamically identifies and optimizes core sub-networks. Through sensitivity-based localization, asynchronous re-selection, and efficient high-rank adaptation, LoSiA achieves high throughput and low activation overhead. Extensive experiments show that LoSiA outperforms baselines on domain-specific and common-sense reasoning tasks while reducing forgetting. We hope that our work will inspire future research to further explore the intrinsic substructures in supervised fine-tuning.
6 Limitation
The innovative design of locating and optimizing sub-networks enables LoSiA to demonstrate outstanding advantages in terms of efficiency and performance. This work preliminarily validates the effectiveness of fine-tuning focused on substructures, yet there remains considerable room for further exploration and improvement. The effectiveness in scenarios such as multi-tasking, vision, and format alignment remains unclear. As for the method, the subnet localization in LoSiA is relatively rigid, and may still fail to precisely capture all critical neuron connections. More flexible and accurate approaches for the location of substructures, such as dynamically adjusting the rank factor for various layers, could further enhance performance.
Furthermore, while LoSiA can be conveniently integrated with other training platforms, additional efforts are required to improve its usability in real-world production scenarios. Currently, our work aims to provide individuals and small enterprises with a highly efficient single-GPU fine-tuning method, but the workflow could be further extended to multi-GPU environments. Moreover, to accommodate diverse datasets and practical deployment conditions, automated time slot selection mechanisms warrant further investigation.
Appendix A Appendix
A.1 Derivations and Proofs
A.1.1 Proof for Formula 1
On a batch $\mathcal{B}$ composed of $M$ samples, the MSE Loss between full fine-tuning (produces model $f$) and training with parameter set $P$ (produces model $f^{\prime}$) is given by:
$\mathcal{L}_{MSE}$ $=\frac{\|y-y^{\prime}\|^{2}_{F}}{M}=\frac{\|Wx-W^{\prime}x\|^{2}_{F}}{M}$ (10)
$\leq\frac{\|W-W^{\prime}\|^{2}_{F}\|x\|^{2}_{F}}{M}$ (11)
SGD
In SGD optimizer, suppose learning rate is $\eta$, the difference in fine-tuned parameter is:
$W-W^{\prime}=-\eta\mathbb{1}_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B})$ (12)
It derives an upper bound for the MSE Loss:
$\mathcal{L}_{MSE}$ $\leq\eta^{2}\frac{\|\mathbb{1}_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B})\|^{2}_{F}\|x\|^{2}_{F}}{M}$ (13)
The result suggests that maximizing the sum of $\nabla_{W_{0}}\mathcal{L}(\mathcal{B})_{ij}$ where $(i,j)\in P$ ideally tightens the approximate error of training on subset.
AdamW
In AdamW optimizer, at training step $t$, the first-order momentum $M_{t}$ and second-order momentum $V_{t}$ are calculated by:
$G_{t}=$ $\nabla_{W}\mathcal{L}(\mathcal{B}_{t})$ (14)
$M_{t}=$ $\beta_{1}M_{t-1}+(1-\beta_{1})G_{t}$ (15)
$V_{t}=$ $\beta_{2}V_{t-1}+(1-\beta_{2})G_{t}^{2}$ (16)
$\tilde{G_{t}}=$ $\frac{M_{t}}{\sqrt{V_{t}+\epsilon}}$ (17)
Similarly, since $W-W^{\prime}=-\eta\mathbb{1}_{(i,j)\not\in P}\cdot\tilde{G_{t}}$, we analysis relationship between $\tilde{G_{t}}$ and $G_{t}$:
$\begin{split}\frac{\partial(\tilde{G_{t}})^{2}}{\partial G_{t}}=&2M_{t}[\frac{(1-\beta_{1})V_{t}}{V_{t}^{2}}\\ &-\frac{(1-\beta_{2})G_{t}M_{t}}{V_{t}^{2}}]\end{split}$ (18)
Suppose $M_{t}>0$, when $G_{t}<\frac{(1-\beta_{1})V_{t}}{(1-\beta_{2})M_{t}}$, $\frac{\partial(\tilde{G_{t}})^{2}}{\partial G_{t}}>0$. In practice, the typical settings are $\beta_{1}=0.9,\beta_{2}=0.999$. Therefore, when $G_{t}<10^{2}\frac{V_{t}}{M_{t}}$, $\tilde{G}_{t}$ increases with $G_{t}$, effectively covering a broad range of non-stationary optimization scenarios.
A.1.2 Proof for Formula 3
The foundational work was established by LeCun et al. (1989) and Kirkpatrick et al. (2016). However, for real-time importance calculation during training, approximations is necessary and is derived bellow. Element-wise importance score $I(\cdot)$ is formulated as:
$\begin{split}I(W_{k})=&|\Delta\mathcal{L}(\mathcal{D})|=|\mathcal{L}(\mathcal{D})-\mathcal{L}_{W_{k}=0}(\mathcal{D})|\\ =&|\frac{\partial\mathcal{L}^{T}(\mathcal{D})}{\partial{W_{k}}}W_{k}-\frac{1}{2}W_{k}^{T}H_{kk}W_{k}\\ &+o(W_{k}^{2})|\end{split}$ (19)
where $H$ denotes the Hessian matrix. It is computational intensive for Hessian calculation, we therefore use fisher information matrix $F$ for diagonal elements instead:
$\begin{split}F_{kk}=&-H_{kk}=-\mathbb{E}_{p(\theta|\mathcal{D})}[\frac{\partial^{2}\mathcal{L}(\theta,D)}{\partial^{2}\theta_{k}}|_{\theta=\theta^{*}}]\\ \approx&-\mathbb{E}_{(x,y)\sim\mathcal{D}}[(\frac{\partial\mathcal{L}(\theta,x,y)}{\partial\theta_{k}}|_{\theta=\theta^{*}})^{2}]\end{split}$ (20)
Approximating the expectation by dataset $\mathcal{D}$ based on Monte Carlo Method, it derives:
$\begin{split}I(W_{k})=&|\frac{\partial\mathcal{L}(\mathcal{D})}{\partial{W_{k}}}W_{k}+o(W_{k}^{2})\\ &-\sum_{(x,y)\in\mathcal{D}}\frac{1}{2|\mathcal{D}|}(\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}})^{2}W_{k}^{2}|\end{split}$ (21)
During training, the dataset $\mathcal{D}$ is processed in batches $\mathcal{B}_{i}$, and the batch gradient is calculated as $\nabla_{W}\mathcal{L}(\mathcal{B}_{i})=\frac{1}{M}\sum_{j=1}^{M}\nabla_{W}\mathcal{L}(\mathcal{B}_{ij})$. To avoid calculating the gradients separately for each sample in the batch, we approximate $\sum_{(x,y)\in\mathcal{B}_{i}}\frac{1}{M}(\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}})^{2}$ to the term $(\frac{\sum_{(x,y)\in\mathcal{B}_{i}}\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}}}{M})^{2}$. To analyze errors, assume $g=\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}}\sim G$, we have:
$\begin{split}\Delta=&|\frac{1}{M}\sum_{j=1}^{M}g_{j}^{2}-(\frac{\sum_{i=1}^{M}{g_{j}}}{M})^{2}|\\ =&\frac{1}{M}\sum_{i=1}^{M}g_{j}^{2}-(\frac{\sum_{j=1}^{M}{g_{j}}}{M})^{2}\\ \leq&\frac{(\max g_{j}-\min g_{j})^{2}}{4}=O(g^{2})\end{split}$ (22)
The approximation errors are bounded. We take the following for importance estimation:
$I_{i}=|\frac{\partial\mathcal{L}(\mathcal{B}_{i})}{\partial{W_{k}}}W_{k}-\frac{1}{2}(\frac{\sum_{j}\frac{\partial\mathcal{L}(\mathcal{B}_{ij})}{\partial{W_{k}}}}{M}W_{k})^{2}\\ +o(W_{k}^{2})|$ (23)
A.1.3 Maximizing Formula 7 is NP-Hard
Task
Given an arbitrary matrix $A^{n\times m}$, select $\tilde{n}\ (\tilde{n}\leq n)$ rows $X_{S}$ and $\tilde{m}\ (\tilde{m}\leq m)$ columns $Y_{S}$, to maximize the sum $\sum_{i\in X_{S}}\sum_{j\in Y_{S}}A_{ij}$.
Lemma
(The Maximum Clique Problem is NP-Complete) Given an undirected graph $G=(V,E)$, where:
  • $V$ is a set of vertices,
  • $E\subseteq V\times V$ is a set of edges,
a clique $C\subseteq V$ is a subset of vertices such that every two distinct vertices in $C$ are adjacent, i.e.,
$\forall u,v\in C\ ,u\neq v\Rightarrow(u,v)\in E.$
The Maximum Clique Problem (MCP) seeks a clique of maximum cardinality in $G$. The problem is NP-complete, meaning:
  • It is NP: a candidate solution can be verified in polynomial time, and
  • It is NP-Hard: any problem in NP can be reduced to it in polynomial time.
Proof
Construct a special form of $A^{n\times n}$ as the adjacent matrix of graph $G$ with larger values on diagonal maximum, that is:
$A_{uv}=\begin{cases}1&\text{if }(u,v)\in E,\\ n^{2}+1&\text{if }u=v\\ 0&\text{otherwise}\end{cases}$
Then, MCP can be reduced to the task in polynomial time following the algorithm:
  • Enumerate $k$ in descending order $n,n-1\ldots 1$
  • Solve the task with $\tilde{n}=\tilde{m}=k$
  • If the optimal solution equals to $(n^{2}+k)k$, then there exist a clique $C=X_{S}$ of size $k$, terminate.
Therefore, a NP-Complete problem can be reduced to the task in polynomial time, which yields the conclusion that the task is NP-Hard.
A.2 Further Observations
A.2.1 Gradient Magnitude Distribution
To investigate the universality of the sparse sub-network structure for large gradients, we analyze gradient magnitude distributions across different layers, as shown in Figure 9. Both Gradients of the Self-Attention and MLP modules exhibits the consistent structure of core subnets.
(a) Decoder Layer $15$
(b) Decoder Layer $25$
Figure 9: Gradient Magnitude Distribution on LLaMA-2 7B for Different Decoder Layers and Modules. Purple curve: row/column gradient sums. Orange curve: smoothed neuron selecting frequency. Best selection strategy for each layers (Row2Column/Column2Row) are record in the title of subplots.
A.2.2 Gradient- or Sensitivity-Based
Figure 10: ARC-E Accuracy under Different Masking Percentage. Linear layers in the 10-th to the 25-th decoder layer of LLaMA-2 7B are masked with gradient-based and sensitivity-based subnet selection strategies.
Figure 10 presents the performance on ARC-E across varying masking percentages. The gradient-based approach identifies the subnet based on magnitude of gradients while masking the remaining parameters. Among importance scoring strategies, the sensitivity-based approach, which is adopted by LoSiA, exhibits stronger robustness in higher masking ratios. However, tuning hyperparameters $\beta_{1},\beta_{2}$ in the EMA of sensitivity-based importance calculation may result in marginal return for LoSiA, as evidenced by the minimal performance gap between the refined and unrefined selection methods.
A.3 Experiments Details
A.3.1 Domain Specific Tasks
We randomly sample 50K data from open-source training datasets: MetaMathQA Yu et al. (2023), Magicoder Wei et al. (2023) and Alpaca-GPT4 Peng et al. (2023), and evaluate fine-tuned models on GSM8K Cobbe et al. (2021), MBPP Austin et al. (2021) and MMLU Hendrycks et al. (2021b, a), respectively. Evaluations are conducted using lm-evaluation-harness Gao et al. (2024), with baseline implementations from LLaMA-Factory Zheng et al. (2024).
Table 6 shows the hyperparameters for fine-tuning LLaMA-2 7B on MetaMathQA. We follow the commonly used configurations for baselines, while aligning GPU memory consumptions. For LoSiA, the hyperparameters for each task and model are listed in Table 7. Rank factor $p$ is set to $\frac{1}{8}$, and the gradient dimension of lm_head is compressed to a fraction by $p_{o}$. Time slot $T$ and learning rate may various across tasks. All experiments are conducted with single run on a NVIDIA A800-80GB GPU and CentOS 7 on x86-64 CPUs. Pytorch version is 2.4.1.
Table 6: Hyperparameter Configurations of Fine-Tuning LLaMA-2 7B on MetaMathQA. Note that $\beta_{1},\beta_{2}$ are EMA smoothing factors in sensitivity-based importance calculation, and are fixed across all experiments. $p$ and $p_{o}$ are dimension factors determining the shape of core subnets. $T$ of LoSiA refers to the time slot between re-selections.
LoRA/DoRA PiSSA GaLore LoSiA
Optimizer AdamW
Epochs 3
Batch Size 4
LR $42e-$ 4 $41e-$ 4 $41e-$ 4 $56e-$ 5
Cutoff Length 2048
Warm-up Ratio 0.1
Rank Related $r=64$ $r=512$ $p=\frac{1}{8},\ p_{o}=\frac{1}{8}$
Scale Related $\alpha=128$ $\alpha=2.0$
Period Related - $T=200$ $T=100$
Others - Full Proj $\beta_{1}=\beta_{2}=0.85$
Implement Layer proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj, lm_head
Table 7: Hyperparameter Configurations of LoSiA across different tasks and models.
Datasets MetaMathQA Magicoder Alpaca-GPT4
LR $56e-$ 5 $55e-$ 5 $55e-$ 5
Time Slot $T$ $100$ $150$ $150$
Rank Factor $p$ $\frac{1}{8}$
Models Gemma-2B LLaMA-2 7B LLaMA-2 13B
Vocabulary Size $256,000$ $32,000$ $32,000$
Dimension Factor $p_{o}$ $\frac{1}{64}$ $\frac{1}{8}$ 1
A.3.2 Common-Sense Reasoning Tasks
Table 8: Datasets of Common-Sense Reasoning.
Datasets #Train #Test Task Type
ARC-CClark et al. (2018) 1,120 1,170 Q & A
ARC-E Clark et al. (2018) 2,250 2,380 Q & A
HellaSwag Zellers et al. (2019) 39,905 10,042 Sentence Completion
Winogrande Sakaguchi et al. (2019) 9,248 1,267 Fill the Blank
PIQA Bisk et al. (2020) 16,100 1,840 Q & A
OBQA Mihaylov et al. (2018) 4,957 500 Q & A
SIQA Sap et al. (2019) 33,410 1,954 Q & A
BoolQ Clark et al. (2019) 9,427 3,270 Text Classification
The datasets of common-sense reasoning tasks are presented in Table 8, while corresponding hyperparameters detailed in Table 9. The GPU memory usage remain aligned. For each PEFT baselines, searches in learning rate are performed.
We report the accuracy metric evaluated by lm-evaluation-harness, which selects answers based on minimal perplexity. This approach mitigates the sensitivity of models to input phrasing variants, thereby enabling a more reliable measurement of the implicit knowledge encoded within the models.
Table 9: Hyperparameter Configurations of Fine-Tuning LLaMA-2 7B on Common-Sense Reasoning Datasets.
LoRA/DoRA PiSSA GaLore LoSiA
Optimizer AdamW
Epochs 3
Batch Size 16
LR $\{1e-4,2e-4\}$$\{5e-5,1e-4\}$$\{1e-4,2e-4\}$$\{5e-5,2e-4\}$
Cutoff Length 256
Warm-up Ratio 0.1
Rank Related $r=64$ $r=512$ $p=\frac{1}{8},\ p_{o}=1$
Scale Related $\alpha=128$ $\alpha=2.0$
Period Related - $T=200$ $T=50$
Others - Full Proj $\beta_{1}=\beta_{2}=0.85$
Implement Layer proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj, lm_head
A.3.3 Rank Factor Robustness
To evaluate the impact of the rank factor $p$, which determines the scale of the core subnets, we conduct an ablation study on MetaMathQA. The results demonstrate LoSiA’s robustness across various subnet scales. Note that $p=1/16$ may be relatively small for effective subnet fine-tuning, while increasing computational budget boosts the performance.
Table 10: Rank Factor Robustness on GSM8K
Model $\mathbf{1/16}$ $\mathbf{1/8}$ $\mathbf{1/4}$ $\mathbf{1/2}$
Gemma-2B $37.53$ $42.84$ $45.03$ $45.64$
LLaMA-2 7B $40.64$ $44.66$ $46.02$ $48.45$
Table 11: The Detail of Ablation Study on MMLU. Note that the variant GL surpasses LoSiA on Humanities but shows performance drop on the rest of domains.
Model MMLU
Humanities Other Social S STEM Avg.
Sensitivity-based Localization (LoSiA) $41.70$ $\mathbf{52.23}$ $\mathbf{50.89}$ $\mathbf{36.82}$ $\mathbf{44.95}$
Gradient-based Localization (GL) $\mathbf{42.64}$ $51.41$ $50.62$ $36.22$ $44.88$
A.3.4 Continue Learning
To examine whether reduction of intruder dimensions in LoSiA mitigates forgetting in continue learning, we sequentially adapt LLaMA-2 7B through five common-sense reasoning tasks by the order HellaSwag, PIQA, BoolQ, SIQA and Winogrande. Learning rate for LoRA is $41e-$ 4 and for LoSiA is $55e-$ 5. The remaining hyperparameters are consistent with Table 9. LoRA modules are merged into the backbone before subsequent task adaptation.
Suppose the model learn sequentially on $N$ tasks. Let $P_{i,j}$ denote the ACC on task $j$ after training on task $i$. Following Zhao et al. (2024b), we formulate the metrics (AP, FWT and BWT) as bellow:
Average Performance: $AP=\frac{1}{N}\sum_{i=1}^{N}P_{N,i}$
Forward Transfer: The metric measures the transferability of learned knowledge from previous tasks to a new task. $\text{FWT}=\frac{1}{N}\sum_{i=1}^{N}(P_{i,i}-P_{0,i})$, where $P_{0,i}$ is the performance of individually training task $i$.
Backward Transfer: The metric evaluates the impact of learning later tasks on the model’s performance on an earlier task, that is $\text{BWT}=\frac{1}{N-1}\sum_{i=1}^{N-1}(P_{N,i}-P_{i,i})$.
Table 12 shows the detailed result during sequential adaptation. After continuing learning through all tasks, Seq-LoSiA outperforms Seq-LoRA across all benchmarks, highlighting its efficiency in forgetting mitigating.
Table 12: Details of Performances on Continue Learning Five Common-Sense Reasoning Tasks. The column stands for training order, while the label "ST" indicates the result in single-tasking training.
Method Task (#1) HellaS (#2) PIQA (#3) BoolQ (#4) SIQA (#5) WinoG ST
Seq-LoRA HellaSwag $59.86$ $55.64$ $59.10$ $57.86$ $54.36$ $59.86$
PIQA $76.01$ $80.52$ $77.86$ $78.73$ $77.64$ $79.33$
BoolQ $77.80$ $73.27$ $86.30$ $80.12$ $75.93$ $88.07$
SIQA $45.80$ $47.80$ $45.85$ $59.52$ $46.11$ $56.86$
Winogrande $64.25$ $68.35$ $68.82$ $69.93$ $79.08$ $73.88$
Seq-LoSiA HellaSwag $63.72$ $61.89$ $61.11$ $60.37$ $\mathbf{56.43}$ $63.72$
PIQA $78.29$ $79.49$ $79.82$ $79.38$ $\mathbf{77.75}$ $81.50$
BoolQ $77.52$ $70.76$ $83.24$ $82.54$ $\mathbf{81.99}$ $84.13$
SIQA $47.80$ $48.26$ $48.26$ $59.93$ $\mathbf{56.04}$ $61.05$
Winogrande $68.51$ $67.88$ $68.51$ $71.82$ $\mathbf{80.19}$ $77.19$
A.4 Resources Measurement
Figure 11 and 12 shows the memory and training time overheads for different PEFT methods on LLaMA-2 7B. With Gradient checkpointing, LoSiA and LoSiA-Pro display lower latency than low-rank methods across all ranks.
Figure 11: GPU Memory Usage and Training Latency Comparison w Gradient Checkpointing
Figure 12: GPU Memory Usage and Training Latency Comparison w/o Gradient Checkpointing
When disables Gradient checkpointing, LoSiA-Pro significantly reduces activation storage by at least $26\%$ and supports $70\%$ additonal training context length under consistent GPU memory constraints compared to LoRA.
A.4.1 Memory Estimate
Consider a model with $L$ decoder layers, each containing $K$ tunable matrices. The model use $b$-bit precision storage, with hidden dimension $d$ and vocabulary size $V$. Table 13 shows GPU memory consumption details of LoRA, GaLore and LoSiA.
Table 13: Comparison of Memory Consumptions. Cells in green highlight the components that may notably lower than other methods, while in red highlight the components that may cause relatively large memory consumption.
LoRA GaLore LoSiA
Update Rank $r$ $R$ $pd$
#Trainable $2LKrdb$ $LKR^{2}b+Vdb$ $LKd^{2}p^{2}b+$$Vdp_{o}b$
#Optimizer $4LKrdb$ $2(LKR^{2}b+Vdb)$ $2(LKd^{2}p^{2}b+$$Vdp_{o}b$)
#Gradient $2LKrdb$ $\max\{d^{2}b,Vdb\}$ $\max\{d^{2}b,Vdb\}$
#Auxiliary $2LKrdb$ $2LKRdb$ $2Kd^{2}b$
#Total $8LKrdb$ $2(LKR^{2}b+Vdb)$ $+\max\{d^{2}b,Vdb\}$ $+2LKRdb$ $2(LKd^{2}p^{2}b+Vdp_{o}b)$ $+\max\{d^{2}b,Vdb\}$ $+2Kd^{2}b$
For optimizers like AdamW, LoSiA reduces the gradient dimension of the output layer to a fraction $p_{o}$, while GaLore performs full fine-tuning on the output layer of shape $d\times V$. Both GaLore and LoSiA utilize per-layer weight update techniques for gradient computation. In contrast, basic LoRA implementations require collecting all gradients after backward propagation. The reduced memory overhead of gradient storage allows GaLore and LoSiA to fully train the $\mathtt{lm\_head}$ in LLaMA-2 models.
In terms of auxiliary parameters, GaLore requires storing down- and up-projection matrices. Since $R$ is typically high-rank, GaLore’s auxiliary parameters can be significantly larger than those of other methods.
For LoSiA, auxiliary parameters are used to compute the importance scores ($\overline{U}(\cdot)$ and $\overline{I}(\cdot)$). If gradient-based importance scoring is adopted, this component can be completely eliminated.
Regarding total memory consumption, increasing the rank in LoRA and GaLore incurs substantial overhead. However, in LoSiA, only the term $2(LKd^{2}p^{2}b+Vdp_{o}b)$ scales with rank factor $p$, resulting in a more efficient memory footprint.
Table 14: Details of Trainable Parameters for LoRA and LoSiA under Different Hyperparameter Configurations on LLaMA-2 7B.
LoRA
$\mathbf{Rank\ r}$ $16$ $64$ $256$ $1024$
#$\mathbf{\#Trainable}$ $40.0$M $160.0$M $640.0$M $2560.0$M
$\mathbf{Mem(GB)}$ $22.33$ $23.72$ $28.42$ $63.90$
LoSiA
$\mathbf{Factor\ p}$ $1/16$ $1/8$ $1/4$ $1/2$
$\mathbf{Update\ Rank\ r}$ $256$ $512$ $1024$ $2048$
$\mathbf{p_{o}=1/8}$
#$\mathbf{\#Trainable}$ $42.8$M $122.1$M $439.3$M $1700.8$M
$\mathbf{Mem(GB)}$ $21.84$ $21.87$ $22.73$ $28.73$
$\mathbf{p_{o}=1}$
#$\mathbf{\#Trainable}$ $158.0$M $238.9$M $562.2$M $1855.7$M
$\mathbf{Mem(GB)}$ $22.24$ $22.84$ $23.37$ $28.98$
A.4.2 Latency Measurement
[…]
Table 15: Comparison of Training Latency on LLaMA-2 7B. Latencies are reported in measuerments of $\mu s$ per token, training with Flash-Attention 2 Dao (2023).
Forward Backward Other Total
w Gradient Check-Pointing
LoRAr=64 74.0 264.0 0 338.0
DoRAr=64 104.2 552.2 0 656.4
GaLoreR=512 70.1 227.5 140.1 (574s / 500 step) 437.7
[…]
w/o Gradient Check-Pointing
LoRAr=64 Out of Memory
[…]
While demonstrating superior performance among existing baselines, LoSiA reduces training latency by 14.1% compared to LoRA, 55.8% compared to DoRA. The acceleration is mainly due to the elimination of low-rank matrix multiplication. Specifically, during backward propagation with Gradient Checkpointing, the production of low-rank matrices introduces significant overhead for activation recomputation and gradient calculation. Note that LoRA can avoid gradient calculations on backbone weights, but this requires specialized implementations and introduces a large coefficient of computational complexity.
For LoSiA-Pro, the computational complexity remains the same as LoSiA during the forward pass, but it only requires storing a proportion $p$ of the input activations of the linear layers. During the backward pass, LoSiA-Pro reduces the computational cost to $p^{2}$ relative to full gradient computation, which significantly lowers the latency of backward propagation. This results in highly efficient training and lower GPU memory consumption.

### v3

@@
 We also propose LoSiA-Pro, a more refined equivalent implementation of LoSiA, which significantly reduces the activation storage and computational complexity in backward propagation.
-LoSiA-Pro speeds up training 1.38×1.38$\times$1.38 × compared to LoRA and 2.68×2.68$\times$2.68 × compared to DoRA.
+LoSiA-Pro speeds up training $1.38\times$ compared to LoRA and $2.68\times$ compared to DoRA.
 In summary, our contributions are as follows.
@@
 (3) We conduct extensive evaluations across multiple models and benchmarks.
-LoSiA outperforms all advanced PEFT baselines on domain-specific and commonsense reasoning tasks, while also accelerating training by 1.15×1.15$\times$1.15 × compared to LoRA.
-Moreover, its efficient variant, LoSiA-Pro, achieves a further speedup of 1.38×1.38$\times$1.38 ×.
+LoSiA outperforms all advanced PEFT baselines on domain-specific and commonsense reasoning tasks, while also accelerating training by $1.15\times$ compared to LoRA.
+Moreover, its efficient variant, LoSiA-Pro, achieves a further speedup of $1.38\times$.
 2 Related Work
@@
 Definition
-[…]
+Consider a model $f_{0}:\mathcal{X}\to\mathcal{Y}$ trained on dataset $\mathcal{D}=\{B_{i}\}_{i=1}^{N}$, where each batch $B_{i}=\{(x_{j},y_{j})\}_{j=1}^{M}$ contains $M$ samples.
+Let $W$ denote the parameters and $\mathcal{L}$ the loss function.
+The neural network $S_{0}$ in $f_{0}$ can be represented as a tuple comprising input neurons $X_{S_{0}}$, output neurons $Y_{S_{0}}$ and neural connections $W_{X_{S_{0}},Y_{S_{0}}}$, that is, $S_{0}=(X_{S_{0}},Y_{S_{0}},W_{X_{S_{0}},Y_{S_{0}}})$.
+The notation $f_{0}\xrightarrow[\mathcal{D}]{P_{0}}f$ denotes training model $f_{0}$ on $\mathcal{D}$ over full parameters $P_{0}=\{W_{X_{S_{0}},Y_{S_{0}}}\}$ and produces model $f$.
+We investigate the following question: Can we efficiently identify a parameter subset $P\subset P_{0}$ of fixed size, such that training $f\xrightarrow[\mathcal{D}]{P}f^{\prime}$ minimizes the loss difference $\Delta\mathcal{L}=|\mathcal{L}(f^{\prime},\mathcal{D})-\mathcal{L}(f,\mathcal{D})|$?
 3.1 Structure of Gradients
@@
 The sparse pattern remains consistent in MLP layers (Appendix A.2.1).
-We therefore focus on fine-tuning subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$ termed the core subnet - rather than the entire network.
+We therefore focus on fine-tuning subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$ - termed the core subnet - rather than the entire network.
 3.2 Subnet Localization
@@
 2:  $\text{sums}\leftarrow\text{Sum}(q,\text{dim}=1)$
-[…]
+3:  $\text{rows}\leftarrow\text{Top-K}(\text{sums},\lfloor np\rfloor).\text{indices}$
 4:  $\text{sums}\leftarrow\text{Sum}(q[\text{rows},:],\text{dim}=0)$
-[…]
-[…]
+5:  $\text{cols}\leftarrow\text{Top-K}(\text{sums},\lfloor mp\rfloor).\text{indices}$
+6:  return $(\text{rows},\text{cols})$
 7:end function
@@
 Consider a model $f$ with $L$ decoder layers $\{D_{l}\}_{l=0}^{L-1}$, where each layer $D_{l}$ contains $K$ linear layers $\{W_{l,k}\}_{k=1}^{K}$ with corresponding core subnets $\{S_{l,k}\}_{k=1}^{K}$.
-The training timeline is divided into time slots of $T$ steps, such that for time slots $[iT,(i+1)T),i=1,2,\ldots$, we:
+The training timeline is divided into time slots of $T$ steps, such that for time slots $iT,(i+1)T),i=1,2,\ldots$, we:
   1. 1.
-    Compute $\overline{I}(\cdot),\overline{U}(\cdot)$ for layer $D_{l}$ in the time slot $[(kL+l-1)T,(kL+l)T),k\in\mathbb{N}$.
+    Compute $\overline{I}(\cdot),\overline{U}(\cdot)$ for layer $D_{l}$ in the time slot $(kL+l-1)T,(kL+l)T),k\in\mathbb{N}$.
   2. 2.
@@
 Formally, the learning rate at step $t$ is:
-$\overline{lr}(t)=\begin{cases}\frac{t-(kL+l)T}{T}\cdot\text{lr}(t)&\text{if C}\\ \text{lr}(t)&\text{otherwise}\\ \end{cases}\$ (8)
-The condition $C$ is $t\in[(kL+l)T,(kL+l+1)T)\text{ and }t>T_{w}$, where $T_{w}$ is the warmup duration.
+[…]
+The condition $C$ is $t\in[(kL+l)T,(kL+l+1)T)\text{ and }t>T_{w}\$, where $T_{w}$ is the warmup duration.
 Figure 4 illustrates the timelines of importance calculation and the learning rate rewarming across layers, with re-localization sandwiched between them.
@@
 5-shot 0-shot,CoT Pass@1 Pass@10 0-shot,PPL 5-shot,GEN
-Gemma 2B FFT $50.1$ $11.0$ $46.4$ $50.4$ $33.0$ $43.4$ $36.1$ $37.0$ $41.05$
+[…]
 LoRA $36.1$ $15.0$ $35.7$ $41.1$ $26.0$ $36.6$ $34.9$ $31.2$ $34.25$
@@
 LoSiA (-Pro) $36.9$ $\mathbf{10.3\ (9.4)}$ $\mathbf{42.8}$ $\mathbf{49.7}$ $30.7$ $43.0$ $\mathbf{37.5}$ $\mathbf{37.4}$ $\mathbf{40.18}$
-LLaMA 2-7B FFT $64.1$ $27.5$ $46.6$ $46.9$ $29.9$ $40.2$ $45.2$ $42.5$ $41.88$
+[…]
 LoRA $23.7$ $33.3$ $42.9$ $46.7$ $26.0$ $37.8$ $42.3$ $37.3$ $38.83$
@@
 LoSiA (-Pro) $21.9$ $\mathbf{26.1\ (21.8)}$ $44.7$ $46.7$ $\mathbf{28.4}$ $\mathbf{39.4}$ $\mathbf{45.0}$ $\mathbf{41.5}$ $\mathbf{40.95}$
-LLaMA 2-13B FFT-8Bit $77.1$ $53.0$ $61.2$ $55.7$ $35.7$ $43.2$ $53.6$ $56.2$ $50.93$
+[…]
 LoRA $36.9$ $56.0$ $58.6$ $\mathbf{56.4}$ $34.1$ $44.8$ $52.6$ $53.7$ $50.03$
@@
 Vanilla LoSiA $\mathbf{44.66}$ $\mathbf{44.95}$ $\mathbf{44.81}$
-Synchronous Localization (SL) $42.76$ $44.13$ $43.45$
-Gradient-based Localization (GL) $43.00$ $44.88$ $43.94$
-w/o Warm-up during Selection (WDS) $38.06$ $44.21$ $41.14$
-w FFT lm_head (FFTO) $43.96$ $44.32$ $44.14$
+Synchronous
+Localization (SL)
+$42.76$ $44.13$ $43.45$
+Gradient-based
+Localization (GL)
+$43.00$ $44.88$ $43.94$
+w/o Warm-up
+during Selection (WDS)
+$38.06$ $44.21$ $41.14$
+w FFT lm_head
+$43.96$ $44.32$ $44.14$
 Figure 6: Loss Curves of Baselines and LoSiA Variants, training on MetaMathQA and Alpaca-GPT4.
@@
 Reduce Intruder Dimensions
-[…]
+Low-rank fine-tuning methods often introduce intruder dimensions Shuttleworth et al.
+(2024b), resulting in spectral discrepancies between the fine-tuned and the pre-trained weights.
+This diminishes the adaptability of LoRA in sequential learning.
+Figure 8 illustrates the cosine similarity between the $\text{Top-}500$ singular vectors of the trained matrices and those of the original weights.
+Both LoRA and DoRA exhibit dimensional shifts due to their low-rank structures, whereas LoSiA demonstrates higher similarity and dimensional stability comparable to FFT.
 Figure 7: Selected Frequency Distributions of Neurons in Core Subnets. Sorted frequencies are ploted in black.
@@
 Method AP($\uparrow$) FWT($\uparrow$) BWT($\uparrow$)
-Seq-LoRA $66.62$ $\mathbf{1.46}$ $-8.04$
-Seq-LoSiA $\mathbf{70.48}$ $-0.20$ $\mathbf{-3.54}$
+Seq-LoRA $66.62$ $\mathbf{1.46}$ 8.04-8.04
+Seq-LoSiA $\mathbf{70.48}$ 0.20-0.20 $\mathbf{-3.54}$
 The results in Table 5 demonstrate that LoSiA outperforms LoRA in mitigating forgetting with 4.5% in BWT and achieves a 3.86% improvement in average performance of sequential fine-tuning.
@@
 Batch Size 4
-LR $42e-$ 4 $41e-$ 4 $41e-$ 4 $56e-$ 5
+LR 2e42e-4 1e41e-4 1e41e-4 6e56e-5
 Cutoff Length 2048
@@
 Rank Related $r=64$ $r=512$ $p=\frac{1}{8},\ p_{o}=\frac{1}{8}$
-Scale Related $\alpha=128$ $\alpha=2.0$
+Scale Related $\alpha=128$ $\alpha=2.0$ -
 Period Related - $T=200$ $T=100$
 Others - Full Proj $\beta_{1}=\beta_{2}=0.85$
-Implement Layer proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj, lm_head
+Implement Layer
+proj_q,proj_k,proj_v,proj_o,
+up_proj,down_proj,gate_proj
+proj_q,proj_k,proj_v,proj_o,
+up_proj,down_proj,gate_proj,
+lm_head
 Table 7: Hyperparameter Configurations of LoSiA across different tasks and models.
 Datasets MetaMathQA Magicoder Alpaca-GPT4
-LR $56e-$ 5 $55e-$ 5 $55e-$ 5
+LR 6e56e-5 5e55e-5 5e55e-5
 Time Slot $T$ $100$ $150$ $150$
@@
 Rank Related $r=64$ $r=512$ $p=\frac{1}{8},\ p_{o}=1$
-Scale Related $\alpha=128$ $\alpha=2.0$
+Scale Related $\alpha=128$ $\alpha=2.0$ -
 Period Related - $T=200$ $T=50$
 Others - Full Proj $\beta_{1}=\beta_{2}=0.85$
-Implement Layer proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj proj_q,proj_k,proj_v,proj_o, up_proj,down_proj,gate_proj, lm_head
+Implement Layer
+proj_q,proj_k,proj_v,proj_o,
+up_proj,down_proj,gate_proj
+proj_q,proj_k,proj_v,proj_o,
+up_proj,down_proj,gate_proj,
+lm_head
 A.3.3 Rank Factor Robustness
@@
 Humanities Other Social S STEM Avg.
-Sensitivity-based Localization (LoSiA) $41.70$ $\mathbf{52.23}$ $\mathbf{50.89}$ $\mathbf{36.82}$ $\mathbf{44.95}$
-Gradient-based Localization (GL) $\mathbf{42.64}$ $51.41$ $50.62$ $36.22$ $44.88$
+Sensitivity-based
+Localization (LoSiA)
+$41.70$ $\mathbf{52.23}$ $\mathbf{50.89}$ $\mathbf{36.82}$ $\mathbf{44.95}$
+Gradient-based
+Localization (GL)
+$\mathbf{42.64}$ $51.41$ $50.62$ $36.22$ $44.88$
 A.3.4 Continue Learning
 To examine whether reduction of intruder dimensions in LoSiA mitigates forgetting in continue learning, we sequentially adapt LLaMA-2 7B through five common-sense reasoning tasks by the order HellaSwag, PIQA, BoolQ, SIQA and Winogrande.
-Learning rate for LoRA is $41e-$ 4 and for LoSiA is $55e-$ 5.
+Learning rate for LoRA is 1e41e-4 and for LoSiA is 5e55e-5.
 The remaining hyperparameters are consistent with Table 9.
@@
 Table 12: Details of Performances on Continue Learning Five Common-Sense Reasoning Tasks. The column stands for training order, while the label "ST" indicates the result in single-tasking training.
-Method Task (#1) HellaS (#2) PIQA (#3) BoolQ (#4) SIQA (#5) WinoG ST
-Seq-LoRA HellaSwag $59.86$ $55.64$ $59.10$ $57.86$ $54.36$ $59.86$
-PIQA $76.01$ $80.52$ $77.86$ $78.73$ $77.64$ $79.33$
-BoolQ $77.80$ $73.27$ $86.30$ $80.12$ $75.93$ $88.07$
-SIQA $45.80$ $47.80$ $45.85$ $59.52$ $46.11$ $56.86$
-Winogrande $64.25$ $68.35$ $68.82$ $69.93$ $79.08$ $73.88$
-Seq-LoSiA HellaSwag $63.72$ $61.89$ $61.11$ $60.37$ $\mathbf{56.43}$ $63.72$
-PIQA $78.29$ $79.49$ $79.82$ $79.38$ $\mathbf{77.75}$ $81.50$
-BoolQ $77.52$ $70.76$ $83.24$ $82.54$ $\mathbf{81.99}$ $84.13$
-SIQA $47.80$ $48.26$ $48.26$ $59.93$ $\mathbf{56.04}$ $61.05$
-Winogrande $68.51$ $67.88$ $68.51$ $71.82$ $\mathbf{80.19}$ $77.19$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 A.4 Resources Measurement
@@
 LoRA GaLore LoSiA
-Update Rank $r$ $R$ $pd$
+Update Rank $r$ $R$ pdpd
 #Trainable $2LKrdb$ $LKR^{2}b+Vdb$ $LKd^{2}p^{2}b+$$Vdp_{o}b$
@@
 #Auxiliary $2LKrdb$ $2LKRdb$ $2Kd^{2}b$
-#Total $8LKrdb$ $2(LKR^{2}b+Vdb)$ $+\max\{d^{2}b,Vdb\}$ $+2LKRdb$ $2(LKd^{2}p^{2}b+Vdp_{o}b)$ $+\max\{d^{2}b,Vdb\}$ $+2Kd^{2}b$
+#Total $8LKrdb$
+$2(LKR^{2}b+Vdb)$
+$\max\{d^{2}b,Vdb\}$
++2LKRdb+2LKRdb
+$2(LKd^{2}p^{2}b+Vdp_{o}b)$
+$\max\{d^{2}b,Vdb\}$
+$2Kd^{2}b$
 For optimizers like AdamW, LoSiA reduces the gradient dimension of the output layer to a fraction $p_{o}$, while GaLore performs full fine-tuning on the output layer of shape $d\times V$.
@@
 $\mathbf{Rank\ r}$ $16$ $64$ $256$ $1024$
-#$\mathbf{\#Trainable}$ $40.0$M $160.0$M $640.0$M $2560.0$M
+$\mathbf{\#Trainable}$ 40.040.0M 160.0160.0M 640.0640.0M 2560.02560.0M
 $\mathbf{Mem(GB)}$ $22.33$ $23.72$ $28.42$ $63.90$
@@
 $\mathbf{p_{o}=1/8}$
-#$\mathbf{\#Trainable}$ $42.8$M $122.1$M $439.3$M $1700.8$M
+$\mathbf{\#Trainable}$ 42.842.8M 122.1122.1M 439.3439.3M 1700.81700.8M
 $\mathbf{Mem(GB)}$ $21.84$ $21.87$ $22.73$ $28.73$
 $\mathbf{p_{o}=1}$
-#$\mathbf{\#Trainable}$ $158.0$M $238.9$M $562.2$M $1855.7$M
+$\mathbf{\#Trainable}$ 158.0158.0M 238.9238.9M 562.2562.2M 1855.71855.7M
 $\mathbf{Mem(GB)}$ $22.24$ $22.84$ $23.37$ $28.98$
 A.4.2 Latency Measurement
-[…]
+We measure the training latency ($\mu s$ / token) fine-tuning with different PEFT methods on LLaMA-2 7B, and the results are shown in Table 15.
+The experiments are conducted with $\text{cutoff\_len}=2048$ and $\text{batch \_size}=4$.
 Table 15: Comparison of Training Latency on LLaMA-2 7B. Latencies are reported in measuerments of $\mu s$ per token, training with Flash-Attention 2 Dao (2023).
@@
 DoRAr=64 104.2 552.2 0 656.4
-GaLoreR=512 70.1 227.5 140.1 (574s / 500 step) 437.7
+GaLoreR=512 70.1 227.5
+140.1
+(574s / 500 step)
+437.7
 […]

### v4

@@
 This design enables effective high-rank adaptation by updating only the sub-network parameters, reducing the additional matrix multiplication.
-We also present LoSiA-Pro, a faster implementation of LoSiA, which reduces the training latency by about $27\%$ compared to LoRA.
+We also present LoSiA-Pro, a faster implementation of LoSiA, which reduces training latency by about $27\%$ compared to LoRA.
 Extensive evaluations show that our method achieves minimal performance drop compared to full fine-tuning, while requiring the least training time across domain specialization and common-sense reasoning tasks.
@@
 1 Introduction
-Figure 1: Overview of LoSiA. It locates and optimizes core sub-network in asynchronous periods.
+Figure 1: Overview of LoSiA. The method locates and optimizes core sub-network in asynchronous periods.
 Large language models, when fine-tuned via supervised learning, can be effectively adapted to downstream tasks such as mathematics Shao et al.
@@
 (2021).
-Although full parameter fine-tuning often yields the best performance, updating billions of parameters is computationally expensive and resource-intensive.
-To address this, parameter-efficient fine-tuning (PEFT) updates only a small subset of parameters to reduce GPU memory usage and communication overhead while maintaining performance comparable to full fine-tuning Houlsby et al.
+Although full parameter fine-tuning often yields the best performance, updating billions of parameters is computationally expensive and resource intensive.
+To address this, parameter-efficient fine-tuning (PEFT) updates only small amount of parameters to reduce GPU memory usage and communication overhead while maintaining performance comparable to full fine-tuning Houlsby et al.
 (2019); Ding et al.
@@
 Among PEFT approaches, LoRA Hu et al.
-(2022) has gained widespread adoption by introducing low-rank matrices to approximate full weight updates, enabling competitive performance with significantly reduced computational and economic costs Taori et al.
+(2022) has gained widespread adoption by introducing low-rank matrices to approximate full weight updates, producing competitive performance with significantly reduced computational and economic costs Taori et al.
 (2023).
@@
 (2024a).
-In such settings, low rank configurations (e.g., $8$ or $16$) can lead to performance degradation and underfitting Biderman et al.
+In such settings, low-rank configurations (e.g., $8$ or $16$) can lead to performance degradation and under-fitting Biderman et al.
 (2024).
@@
 Recent studies have attempted to approximate high-rank updates by accumulating multiple low-rank components.
-However, these approaches still suffer from issues such as locally low-rank updates Meng et al.
-(2024b); Lialin et al.
-(2023) or increased computational complexity Zhao et al.
+However, these approaches still suffer from issues such as locally low-rank updates Lialin et al.
+(2023); Meng et al.
+(2024b) or increased computational complexity Zhao et al.
 (2024a).
-Therefore, while the low-rank assumption offers notable improvements in efficiency, it also introduces inherent limitations.
+Therefore, while the low-rank assumption offers notable improvements in parameter efficiency, it also introduces inherent limitations.
 The Lottery Ticket Hypothesis Frankle and Carbin (2019) suggests that dense neural networks contain trainable sub-networks capable of achieving comparable test accuracy.
-This prompts us to reconsider the route of PEFT and explore an alternative: Can we identify and fine-tune such sub-networks within the backbone model to achieve high-quality adaptation more efficiently?
+This prompts us to reconsider the PEFT roadmap and explore an alternative: Can we identify and fine-tune such sub-networks within the backbone model to achieve high-quality adaptation more efficiently?
 To answer this question, we propose LoSiA (Low-Resources Subnet Integration Adaptation), a novel PEFT framework that dynamically localizes and optimizes critical sub-networks periodically, as illustrated in Figure 1.
@@
 Following localization, it fine-tunes the identified sub-network and applies a rewarming learning rate strategy to promote stable and robust training.
-The design enables real-time high-rank updates without introducing additional matrix multiplication overhead, which reduces training latency and ensures no steep increase in training time for high-rank updating.
-Additionally, LoSiA does not introduce extra architectural components and only requires an optimizer replacement for seamless deployment.
+The design enables real-time high-rank updates without introducing additional matrix multiplication overhead, while significantly reducing training latency.
+Additionally, LoSiA does not introduce extra architectural components and only requires optimizer replacements for seamless deployment.
 Extensive experiments demonstrate its superior performance among PEFT baselines on domain-specific, commonsense reasoning tasks, while mitigating forgetting in continue learning.
@@
 In summary, our contributions are as follows.
-(1) Innovatively introduces the structure of the sub-network to the field of parameter-efficient fine-tuning.
-We develop periodic subnet localization, optimization and integration techniques to dynamically capture and adapt task-essential parameters.
-(2) We propose LoSiA, a novel PEFT approach that dynamically localizes and optimizes sub-networks to achieve high performance, efficiency, and usability.
-To further improve practicality, we also introduce a faster variant, LoSiA-Pro.
+(1) Innovatively, we incorporate sub-network structure into the field of parameter-efficient fine-tuning.
+We devise a periodic workflow with techniques that localize, optimize, and integrate sub-networks, thus flexibly capturing and adapting task-essential parameters.
+(2) We propose LoSiA, a novel high performance PEFT approach that dynamically localizes and optimizes sub-networks.
+By eliminating redundant computation, we further propose LoSiA-Pro, a loss-less variant that markedly reduces training latency and GPU memory footprint.
 (3) We conduct extensive evaluations across multiple models and benchmarks.
-LoSiA outperforms all advanced PEFT baselines on domain-specific and commonsense reasoning tasks, while also accelerating training by $1.15\times$ compared to LoRA.
+LoSiA outperforms all advanced PEFT baselines on domain-specific and common-sense reasoning tasks, while also accelerating training $1.15\times$ compared to LoRA.
 Moreover, its efficient variant, LoSiA-Pro, achieves a further speedup of $1.38\times$.
@@
 Parameter-Efficient Fine-Tuning
-Full parameter fine-tuning (FFT) adapts pre-trained models to downstream tasks by updating all model parameters Wei et al.
-(2022), but it is often computationally expensive.
-In contrast, parameter-efficient fine-tuning (PEFT) methods update only a small subset of parameters, significantly reducing training costs while still maintaining strong performance.
+Full param-eter fine-tuning (FFT) adapts pre-trained models to downstream tasks by updating all model parameters Wei et al.
+(2022), yet incurs prohibitive computational overhead.
+In contrast, parameter-efficient fine-tuning (PEFT) methods update only a small subset of parameters, curbing training costs while sustaining competitive accuracy.
 LoRA Hu et al.
-(2022) approximates parameter updates as the product of low-rank matrices, achieving promising performance for tasks such as instruction tuning Ghosh et al.
+(2022) approximates parameter updates as the product of low-rank matrices, achieving promising performance in tasks such as instruction tuning Ghosh et al.
 (2024).
 Enhanced variants such as PiSSA Meng et al.
-(2024a) accelerate convergence by prioritizing major singular vectors, while DoRA Liu et al.
-(2024) improves performance in low-rank by decomposing updates into directional and magnitude components.
+(2024a) accelerate convergence by prioritizing dominant singular vectors, while DoRA Liu et al.
+(2024) decomposes updates into directional and magnitude components for more effective fine-tuning.
 Other derivatives such as LoRA+ Hayou et al.
 (2024b), LoRA-GA Wang et al.
-(2024a) and LoRA-Dash Si et al.
-(2025) refine the framework by directional or module biased fine-tuning.
+(2024a), and LoRA-Dash Si et al.
+(2025) refine the framework by directional or module biased optimization.
 However, recent studies Jiang et al.
@@
 (2024); Ghosh et al.
-(2024) reveal that the low-rank structure limits effectiveness in knowledge-intensive domains (e.g., mathematics, coding).
-Advanced solutions adopt novel strategies: 1) Architectural modifications through MoE-based LoRA combinations Zadouri et al.
+(2024) reveal that the low-rank bottleneck restricts effectiveness in knowledge-intensive domains (e.g., mathematics, coding).
+Advanced solutions adopt strategies such as: 1) Architectural modifications through MoE-based LoRA combinations Zadouri et al.
 (2023); Li et al.
 (2024); Wang et al.
-(2024b) for multitask scenarios; 2) High-rank fine-tuning via accumulated low-rank projections, such as ReLoRA Lialin et al.
+(2024b) for multitask learning scenarios; 2) High-rank fine-tuning via accumulated low-rank projections, such as ReLoRA Lialin et al.
 (2023), MoRA Jiang et al.
@@
 (2024a) to enhance training effectiveness.
-However, these approaches incur an increase in architectural complexity or a drop in throughput.
-Rare methods achieve an optimal balance between performance, training latency, and implementation simplicity.
+However, these ameliorated approaches either inflate architectural complexity or compromise throughput.
+Rare methods simultaneously optimize performance, training latency, and implementation simplicity.
 Skill Localization and Pruning
-LLM pruning reduces neural network size by eliminating redundant or less critical parameters.
+LLM pruning compresses networks by excising redundant or less critical parameters.
 Previous work demonstrates that sparse networks can play crucial roles Frankle and Carbin (2019); Yao et al.
@@
 Panigrahi et al.
-(2023) identifies critical parameters in fine-tuned LMs by optimizing masks of grafted models, although such methods require additional training time and data.
+(2023) identifies critical parameters in fine-tuned LMs by optimizing masks of grafted models, but such methods require additional training time and data.
 Alternatively, gradient- and sensitivity-based metrics enable real-time identification of task-aware parameters Molchanov et al.
@@
 (2022).
-Recent advances adapt the techniques to PEFT: Zhang et al.
-(2023) prunes LoRA trainable parameters, while Feng et al.
-(2024) applies the approach to continual learning scenarios.
+Recent advances extend these ideas to PEFT: Zhang et al.
+(2023) prunes LoRA trainable parameters, while KIF (Feng et al., 2024a, b) integrates skill localization into continual-learning regimes.
 3 Method
@@
 Let $W$ denote the parameters and $\mathcal{L}$ the loss function.
-The neural network $S_{0}$ in $f_{0}$ can be represented as a tuple comprising input neurons $X_{S_{0}}$, output neurons $Y_{S_{0}}$ and neural connections $W_{X_{S_{0}},Y_{S_{0}}}$, that is, $S_{0}=(X_{S_{0}},Y_{S_{0}},W_{X_{S_{0}},Y_{S_{0}}})$.
-The notation $f_{0}\xrightarrow[\mathcal{D}]{P_{0}}f$ denotes training model $f_{0}$ on $\mathcal{D}$ over full parameters $P_{0}=\{W_{X_{S_{0}},Y_{S_{0}}}\}$ and produces model $f$.
-We investigate the following question: Can we efficiently identify a parameter subset $P\subset P_{0}$ of fixed size, such that training $f\xrightarrow[\mathcal{D}]{P}f^{\prime}$ minimizes the loss difference $\Delta\mathcal{L}=|\mathcal{L}(f^{\prime},\mathcal{D})-\mathcal{L}(f,\mathcal{D})|$?
+The neural sub-network $S$ in $f_{0}$ is represented as the tuple $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$, comprising its input neurons $X_{S}$, output neurons $Y_{S}$ and neural connections $W_{X,Y}$.
+Training model $f_{0}$ on $\mathcal{D}$ with full parameters $P_{0}=W$ is compactly written as $f_{0}\xrightarrow[\mathcal{D}]{P_{0}}f$.
+We investigate the following question: Given a cardinality budget, can we efficiently identify a parameter subset $P\subset P_{0}$, such that $f_{0}\xrightarrow[\mathcal{D}]{P}f^{\prime}$ minimizes the loss difference $\Delta\mathcal{L}=|\mathcal{L}(f^{\prime},\mathcal{D})-\mathcal{L}(f,\mathcal{D})|$?
 3.1 Structure of Gradients
-Inspired by pruning techniques, we minimize the mean squared error (MSE) $\mathcal{L}_{\text{MSE}}$ between the outputs of models $f_{k},f^{\prime}_{k}$, where both are trained on $f_{k-1}$ with trainable parameter set $P_{0}$ and $P$, respectively.
+Inspired by pruning techniques, we expect to minimize the mean squared error (MSE) $\mathcal{L}_{\text{MSE}}$ between subsequent models $f_{k}$ and $f^{\prime}_{k}$, where they both are trained from the model $f_{k-1}$ with trainable parameters $P_{0}$ and $P$, respectively.
 For SGD optimizers, we derive the bound:
 $\scriptstyle\mathcal{L}_{MSE}$ $\leq\eta^{2}\frac{\|1_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B}_{k})\|^{2}_{F}\|x\|^{2}_{F}}{M}$ (1)
-For AdamW optimizers, $\mathcal{L}_{MSE}$ admits a similar bound through $\nabla W$ in most cases (Appendix A.1.1).
-Thus, gradient magnitudes in $P$ provide an upper bound for the approximation error, while prioritizing larger gradients yields tighter bounds.
-We therefore seek optimal subsets $P$ to capture the parameters with large gradient magnitudes.
-Ideally, selecting the $top\text{-}K$ entries of $\nabla W$ would suffice, but storing and fine-tuning sparse matrices compromises efficiency.
-Instead, we claim that a suitable pattern for $P$ corresponds to the parameters in subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$, i.e., all connections between the input neuron set $X_{S}$ and the output neuron set $Y_{S}$.
+For AdamW, an analogous bound $\mathcal{L}_{MSE}$ holds in terms of $\nabla W$ in most cases (Appendix A.1.1).
+Thus, gradient magnitudes lay in $P$ provide an upper bound for the approximation error, while retaining parameters with the largest gradient magnitudes to adjust tightens the bound.
+Ideally, selecting the $\text{Top-}K$ entries of $\nabla W$ is theoretically optimal, but storing and fine-tuning sparse matrices compromises the efficiency.
+Instead, we claim that a suitable selection pattern for $P$ corresponds to a structured subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$, i.e., all connections between input neuron set $X_{S}$ and output neuron set $Y_{S}$.
 Figure 2: Gradient Magnitude Distribution of proj_v. Large gradients follow a sparse subnet distribution.
-To validate this selection paradigm, Figure 2 visualizes the gradient magnitude distributions in LLaMA-2 7B’s proj_v layer: The 32 attention heads exhibit a significant disparity in attention scores, while gradients strongly correlate with output neurons $Y_{S}$.
-In particular, a consistent subset of input neurons $X_{S}$ (green markers, x-axis) contributes dominantly to all attention heads.
-The sparse pattern remains consistent in MLP layers (Appendix A.2.1).
-We therefore focus on fine-tuning subnet $S=(X_{S},Y_{S},W_{X_{S},Y_{S}})$ - termed the core subnet - rather than the entire network.
+To validate this selection paradigm, Figure 2 visualizes the gradient magnitude distributions in LLaMA-2 7B’s proj_v layer.
+Across the 32 attention heads, the gradient norms exhibit pronounced skewness and are highly correlated with the corresponding output neurons $Y_{S}$.
+On the other hand, a consistent set of input neurons $X_{S}$ (green markers, x-axis) contributes dominantly to all attention heads.
+The sparse pattern also holds in MLP layers (Appendix A.2.1).
+Consequently, we restrict the fine-tuning space to subnet structures $S$ - termed the core subnet - rather than the entire network.
 3.2 Subnet Localization
-To localize core subnets efficiently, an ideal algorithm should satisfy three key requirements: 1) Efficiency: no additional data requirements or significant latency.
-2) Lightweight: minimal GPU memory overhead.
-3) Dynamic Awareness: enable real-time localizations during training.
-To meet these objectives, the subnet localization process is divided into two stages:
+To efficiently localize core subnets, an ideal algorithm should satisfy three key requirements: 1) Efficiency: no extra data or heavy computation.
+2) Lightweight: negligible GPU memory overhead.
+3) Dynamic Awareness: enable on-the-fly localization throughout the training process.
+Although existing LLM-pruning methods have achieved impressive compression ratios, they still fall short of simultaneously satisfying the aforementioned desiderata.
+Consequently, we devise a dedicated subnet-localization algorithm tailored for efficient fine-tuning that is divided into two stages:
 Parameter Importance Calculation
@@
 (1989); Ma et al.
-(2023) observe the change in loss assuming $W_{k}=0$ for the $k$-th parameter.
-Adopting second-order Taylor expansion, element-wise $I(\cdot)$ is estimated as:
+(2023) observe the change in loss by assuming $W_{k}=0$ for the $k$-th parameter.
+Adopting the second-order Taylor expansion, element-wise importance is estimated as:
 […]
@@
 […]
-Furthermore, estimates by single micro-batch may introduce bias by ignoring training dynamics.
+Furthermore, estimation with single micro-batch may inject bias by overlooking training dynamics.
 Sensitivity smoothing and uncertainty quantification Zhang et al.
 (2022) are used to handle the problem.
-For the training step $i$, compute an exponential moving average (EMA) $\overline{I}_{i}$ for $I_{i}$ ,and uncertainty $\overline{U}_{i}$ for variation $\Delta I_{i}=|I_{i}-\overline{I}_{i}|$:
+At training step $i$, it maintains an exponential moving average (EMA) $\overline{I}_{i}$ for $I_{i}$, and uncertainty $\overline{U}_{i}$ for variation $\Delta I_{i}=I_{i}-\overline{I}_{i}$:
 $\overline{I}_{i}(W_{k})$ $=\beta_{1}\overline{I}_{i-1}(W_{k})+(1-\beta_{1})I_{i}(W_{k})$ (4)
@@
 $s(W_{k})$ $=\overline{I}(W_{k})\cdot\overline{U}(W_{k})$ (6)
-where $\beta_{1},\beta_{2}\in(0,1)$ is the EMA factors.
-We regard $s(\cdot)$ as a appropriate importance assessment.
-Notably, to obtain gradients $W$ asynchronously, LoSiA use per-layer weight updates Lv et al.
-(2024), executing the optimization during backpropagation without storing gradients.
+where $\beta_{1},\beta_{2}\in(0,1)$ are the EMA factors.
+We treat $s(\cdot)$ as an appropriate importance assessment.
+To obtain the weight-gradient signal without keeping all full tensors in memory, LoSiA uses per-layer updates Lv et al.
+(2024), executing optimizations during backpropagation without storing gradients.
 Core Subnet Localization via Importance Scores
-For a subnet $S$ selected from origin network $S_{0}=(\{i\}_{i=1}^{n},\{j\}_{j=1}^{m},W)$, define its importance as:
+Given a subnet $S$ of the origin network $S_{0}=(\{i\}_{i=1}^{n},\{j\}_{j=1}^{m},W)$, define its importance as:
 $s(S)=\sum_{i\in X_{S}}\sum_{j\in{Y_{S}}}s(W_{ij})$ (7)
-The objective is to identify optimal subset $S$ and maximize $s(S)$, while satisfying the memory constraint $\max\{\frac{|X_{S}|}{n},\frac{|Y_{S}|}{m}\}\leq p$, where $p\in(0,1)$ represents the rank factor.
+Our objective is to identify the optimal subnet $S$ that maximizes $s(S)$, while respecting the memory cap $\max\{\frac{|X_{S}|}{n},\frac{|Y_{S}|}{m}\}\leq p$, where $p\in(0,1]$ represents the rank factor.
 However, the task is NP-Hard.
-Leveraging observations in Section 3.1 regarding gradient magnitude sparsity patterns, we develop greedy selection algorithms to identify critical input and output neuron sets $(X_{S},Y_{S})$:
+Exploiting the gradient-magnitude sparsity patterns observed in Section 3.1, we develop greedy selection algorithms to select the critical input and output neuron set $X_{S}$ and $Y_{S}$.
 Algorithm 1 Greedy Strategy for Localization
-1:function Row2Column($q$)
-2:  $\text{sums}\leftarrow\text{Sum}(q,\text{dim}=1)$
-3:  $\text{rows}\leftarrow\text{Top-K}(\text{sums},\lfloor np\rfloor).\text{indices}$
-4:  $\text{sums}\leftarrow\text{Sum}(q[\text{rows},:],\text{dim}=0)$
-5:  $\text{cols}\leftarrow\text{Top-K}(\text{sums},\lfloor mp\rfloor).\text{indices}$
-6:  return $(\text{rows},\text{cols})$
+1:Importance matrix $s\in\mathbb{R}^{n\times m}$ with $s_{ij}=s(W_{ij})$; rank factor $p\in(0,1]$.
+[…]
+3:function Row2Column($s\in\mathbb{R}^{n\times m},\ p$)
+4:  $\rho\leftarrow\operatorname{Top-K\ Indices}\!\ \bigl(\sum_{j=1}^{m}s_{:,j},\;\lfloor np\rfloor\bigr)$
+5:  $\gamma\leftarrow\operatorname{Top-K\ Indices}\!\ \bigl(\sum_{i\in\rho}s_{i,:},\;\lfloor mp\rfloor\bigr)$
+6:  return $(\rho,\gamma)$
 7:end function
-Algorithm 1 implements a row-major greedy approach that selects the $top\text{-}K$ rows by importance score summation, and then finds $\{Y_{S}\}$ to maximum $s(S)$ given fixed $\{X_{S}\}$.
-A column-major greedy algorithm Column2Row is also considered.
-The better selection result of two strategies is selected.
+Algorithm 1 embodies a row-major greedy policy.
+First, it locks the $\lfloor np\rfloor$ input neurons with the highest row-wise aggregate importance, then greedily retains the $\lfloor mp\rfloor$ output neurons that maximize the residual mass in those fixed rows.
+A symmetric column-major variant reverses the order of fixation.
+The final subnet adopts whichever of the two masks yields the higher score $s(S)$.
 Dimensionality Reduction in Output Layer Fine-Tuning
-While prior work Chen et al.
-(2024) has established the benefits of fine-tuning the output layer in conjunction with PEFT methods, the approach remains computationally prohibitive for large-vocabulary models (e.g., Gemma-2B).
+Although prior work Chen et al.
+(2024) has established the benefits of fine-tuning the output layer in conjunction with PEFT methods, the approach remains computationally prohibitive for large-vocabulary models (e.g.
+Gemma-2B).
 However, empirically, backward propagation through the output layer exhibits gradient sparsity, with only a limited subset of tokens receiving significant updates.
-Building on this insight, LoSiA easily implements an efficient optimization strategy by constructing a tunable subnet $S=(X_{S_{0}},Y_{S},W_{X_{S_{0}},Y_{S}})$ of the output layer, where $|Y_{S}|=p_{o}|Y_{S_{0}}|$, and $p_{o}\in(0,1)$ denoting the dimension reduction factor.
+Building on this insight, LoSiA easily implements an efficient optimization strategy by constructing a tunable subnet $S=(X_{S_{0}},Y_{S},W_{X_{S_{0}},Y_{S}})$ in the output layer, where $|Y_{S}|=p_{o}|Y_{S_{0}}|$, and $p_{o}\in(0,1]$ denote the output dimension reduction factor.
 3.3 Subnet Optimization and Intergration
-During fine-tuning, the locations of core subnets may undergo dynamic shifts, as illustrated in Figure 3.
+During fine-tuning, the locations of core subnets undergo dynamic shifts, as illustrated in Figure 3.
+Figure 3: Core Subnet Distribution during Training. The chosen subnets various across training iterations.
 Although a small subset of neurons is consistently selected, peripheral components exhibit significant temporal variability.
-Fine-tuning with a static subnet risks model underfitting and neuron over-specialization.
-To address the issue, we introduce an asynchronous and periodic subnet re-localization mechanism that adapts to the evolving network topology.
-Figure 3: Core Subnet Distribution during Training. The subnet various across different training iterations.
-Naive periodic re-localization strategies can induce training instability and loss spikes Lialin et al.
+Freezing a fixed mask therefore invites under-fitting and over-specialization of the lucky prophase winners.
+To address the issue, we introduce an asynchronous periodic subnet re-localization mechanism that adapts to the evolving network topology.
+Naive periodic learning strategies can induce training instability and loss spikes Lialin et al.
 (2023).
-Furthermore, the storage requirements for $\overline{I}(\cdot),\overline{U}(\cdot)$ in a synchronous period would lead to a scaling of GPU memory overhead.
-Therefore, we propose asynchronous periodic localization coupled with rewarmings of learning rate.
-Consider a model $f$ with $L$ decoder layers $\{D_{l}\}_{l=0}^{L-1}$, where each layer $D_{l}$ contains $K$ linear layers $\{W_{l,k}\}_{k=1}^{K}$ with corresponding core subnets $\{S_{l,k}\}_{k=1}^{K}$.
-The training timeline is divided into time slots of $T$ steps, such that for time slots $iT,(i+1)T),i=1,2,\ldots$, we:
+Furthermore, the storage requirements of $\overline{I}(\cdot),\overline{U}(\cdot)$ for every layer simultaneously would lead to a scaling of GPU memory overhead.
+Therefore, we propose asynchronous periodic localization coupled with rewarmings of learning rate techniques.
+Consider a model $f$ with $L$ decoder layers $\{D_{l}\}_{l=0}^{L-1}$, where each decoder $D_{l}$ contains $K$ linear layers $\{W_{l,k}\}_{k=1}^{K}$, with corresponding core subnets $\{S_{l,k}\}_{k=1}^{K}$.
+The training timeline is chopped into time slots of length $T$, such that for time slots $iT,(i+1)T),i=1,2,\ldots$, we:
   1. 1.
-    Compute $\overline{I}(\cdot),\overline{U}(\cdot)$ for layer $D_{l}$ in the time slot $(kL+l-1)T,(kL+l)T),k\in\mathbb{N}$.
+    Calculate $\overline{I}(\cdot),\overline{U}(\cdot)$ for layer $D_{l}$ in time slots $(kL+l-1)T,(kL+l)T),k\in\mathbb{N}$.
   2. 2.
-    Sequentially reselect $S_{l}$ by $s(\cdot)$ before step $t=(kL+l)T$, the end of time slots in 1.
-This yields a reselection period of $\overline{T}=LT$ for each layer.
-The asynchronous design ensures that at any step, there is exactly one decoder layer calculating $\overline{I}(\cdot),\overline{U}(\cdot)$ and one rewarming, which greatly reduces additional GPU memory overheads for importance score calculation.
-We also implement a learning rate rewarm-up to further enhance training stability.
-Formally, the learning rate at step $t$ is:
+    Sequentially reselect $S_{l}$ by $s(\cdot)$ before step $t=(kL+l)T$, the end of time slots.
+Consequently, every core subnet is refreshed exactly once every $\overline{T}=LT$ steps, and, at any moment, only one layer is (i) accumulating importance statistics and (ii) rewarming learning rate.
+This greatly reduces the extra GPU memory footprint for importance score calculation.
+The rewarming mechanism resets the learning rate to a short warm-up schedule to enhance training stability. Formally, the learning rate at step $t$ is:
 […]
-The condition $C$ is $t\in[(kL+l)T,(kL+l+1)T)\text{ and }t>T_{w}\$, where $T_{w}$ is the warmup duration.
-Figure 4 illustrates the timelines of importance calculation and the learning rate rewarming across layers, with re-localization sandwiched between them.
+The condition $Cond$ is $t\in[(kL+l)T,(kL+l+1)T)\text{ and }t>T_{w}\$, where $T_{w}$ is the warmup duration.
+This means that rewarmings are triggered only after the initial warmup phase is finished.
+Figure 4 illustrates the timelines of importance calculation and the rewarming procedure across multiple layers.
+Importance scores are evaluated, immediately followed by learning rate rewarming, with re-localization sandwiched between them.
 Figure 4: Asynchronous Periodic Subnet Reselection and Learning Rate Rewarming Mechanism (in a 5-layer model for example).
 3.3.1 Faster Implementation (LoSiA-Pro)
-Through subnet fine-tuning, LoSiA can further mitigate activation storage and backward latency. The gradient for the subnet $S$ can be factorized as:
+Through subnet fine-tuning, LoSiA can further mitigate activation storage and backward latency. The gradient of subnet $S$ can be factorized as:
 $\begin{split}\frac{\partial\mathcal{L}}{\partial W_{S}}&=\frac{\partial\mathcal{L}}{\partial W}[X_{S},:][:,Y_{S}]\\ &=(x^{T}[X_{S},:])(\frac{\partial\mathcal{L}}{\partial y}[:,Y_{S}])=\tilde{L}_{S}\tilde{R}_{S}\end{split}$ (9)
-Noticing $\tilde{L}_{S}\in\mathbb{R^{np\times bs}},\tilde{R}_{S}\in\mathbb{R}^{bs\times mp}$, the input activation storage is reduced by a factor $p$, while the computational complexity of gradient calculation is reduced from $O(nmbs)$ to $O(nmbsp^{2})$.
+Noticing $\tilde{L}_{S}\in\mathbb{R}^{np\times bs},\tilde{R}_{S}\in\mathbb{R}^{bs\times mp}$, the input activation storage is reduced by a factor $p$, while the computational complexity of the gradient calculation is reduced from $O(nmbs)$ to $O(nmbsp^{2})$.
 We named the method LoSiA-Pro, a refined equivalent implementation of LoSiA.
-It offers a 27.6% latency reduction compared to LoRA with Gradient Check-Pointing, while reducing 13.4GB GPU memory consumption compared to LoSiA training without Gradient Check-Pointing.
+It offers a 27.6% latency reduction compared to LoRA, while additionally reducing 13.4GB GPU memory consumption compared to LoSiA when training without Gradient Check-Pointing.
 4 Experiments
-We evaluate LoSiA across a broad range of model scales and datasets, conducting rigorous comparisons with common baselines.
+We evaluate LoSiA across a wide range of model scales and datasets, conducting rigorous comparisons with common baselines.
 On both domain-specific and common-sense reasoning tasks, the method demonstrates robust performance with significantly reduced training overheads.
@@
 Specifically, training sets are sampled by 50,000 random entries from MetaMathQA, Magicoder, and Alpaca-GPT4, respectively.
-The GSM8K, MBPP, and MMLU benchmark are for testing.
+The GSM8K, MBPP, and MMLU benchmarks are for testing.
 Additionally, we also compared LoSiA with baseline methods on eight common sense reasoning tasks.
@@
 Table 1: Comparison of PEFT Methods Across Models on Domain-Specific Tasks.
-Accuracy is reported, alongside with memory consumption (GB), maximum per-task training time (h).
-The numbers in parentheses indicate the training time of LoSiA-Pro, which is a refined and computationally equivalent implementation of LoSiA.
-Model Method Mem(GB) Time(h) GSM8K MBPP MMLU Avg.
+Accuracy is reported, alongside with memory consumption (GB) and per-token training latency ($\mu s$ / token).
+The numbers in parentheses indicate the latency of LoSiA-Pro, which is a refined and computationally equivalent implementation of LoSiA.
+Model Method Mem
+Latency
+($\mu s$ / token)
+GSM8K MBPP MMLU Avg.
 5-shot 0-shot,CoT Pass@1 Pass@10 0-shot,PPL 5-shot,GEN
-[…]
-LoRA $36.1$ $15.0$ $35.7$ $41.1$ $26.0$ $36.6$ $34.9$ $31.2$ $34.25$
-PiSSA $36.1$ $14.9$ $38.5$ $46.5$ $26.4$ $39.0$ $33.8$ $32.6$ $36.13$
-DoRA $37.3$ $29.8$ $39.7$ $43.0$ $31.4$ $\mathbf{43.2}$ $36.2$ $37.1$ $38.43$
-GaLore $37.5$ $14.2$ $39.3$ $44.7$ $\mathbf{31.6}$ $42.6$ $36.6$ $35.5$ $38.38$
-LoSiA (-Pro) $36.9$ $\mathbf{10.3\ (9.4)}$ $\mathbf{42.8}$ $\mathbf{49.7}$ $30.7$ $43.0$ $\mathbf{37.5}$ $\mathbf{37.4}$ $\mathbf{40.18}$
-[…]
-LoRA $23.7$ $33.3$ $42.9$ $46.7$ $26.0$ $37.8$ $42.3$ $37.3$ $38.83$
-PiSSA $23.7$ $33.1$ $43.5$ $46.2$ $26.8$ $36.6$ $42.7$ $38.5$ $39.05$
-DoRA $24.2$ $68.3$ $\mathbf{45.0}$ $\mathbf{47.2}$ $26.0$ $34.4$ $44.1$ $36.7$ $38.90$
-GaLore $23.7$ $39.6$ $42.2$ $45.3$ $28.0$ $39.0$ $43.1$ $41.2$ $39.80$
-LoSiA (-Pro) $21.9$ $\mathbf{26.1\ (21.8)}$ $44.7$ $46.7$ $\mathbf{28.4}$ $\mathbf{39.4}$ $\mathbf{45.0}$ $\mathbf{41.5}$ $\mathbf{40.95}$
-[…]
-LoRA $36.9$ $56.0$ $58.6$ $\mathbf{56.4}$ $34.1$ $44.8$ $52.6$ $53.7$ $50.03$
-PiSSA $36.9$ $55.5$ $53.4$ $55.2$ $34.5$ $44.8$ $52.0$ $48.8$ $48.11$
-LoSiA (-Pro) $36.9$ $\mathbf{46.5\ (38.6)}$ $\mathbf{59.0}$ $54.0$ $\mathbf{34.9}$ $\mathbf{48.2}$ $\mathbf{53.1}$ $\mathbf{55.7}$ $\mathbf{50.82}$
+Gemma 2B FFT $50.1$ $142.9$ $46.4$ $50.4$ $33.0$ $43.4$ $36.1$ $37.0$ $41.05$
+LoRA $36.1$ $136.7$ $35.7$ $41.1$ $26.0$ $36.6$ $34.9$ $31.2$ $34.25$
+PiSSA $36.1$ $136.9$ $38.5$ $46.5$ $26.4$ $39.0$ $33.8$ $32.6$ $36.13$
+DoRA $37.3$ $296.6$ $39.7$ $43.0$ $31.4$ $\mathbf{43.2}$ $36.2$ $37.1$ $38.43$
+GaLore $37.5$ $162.4$ $39.3$ $44.7$ $\mathbf{31.6}$ $42.6$ $36.6$ $35.5$ $38.38$
+LoSiA (-Pro) $36.9$ $\mathbf{119.9\ (107.2)}$ $\mathbf{42.8}$ $\mathbf{49.7}$ $30.7$ $43.0$ $\mathbf{37.5}$ $\mathbf{37.4}$ $\mathbf{40.18}$
+LLaMA 2-7B FFT $64.1$ 359.2 $46.6$ $46.9$ $29.9$ $40.2$ $45.2$ $42.5$ $41.88$
+LoRA $23.7$ $338.0$ $42.9$ $46.7$ $26.0$ $37.8$ $42.3$ $37.3$ $38.83$
+PiSSA $23.7$ $338.5$ $43.5$ $46.2$ $26.8$ $36.6$ $42.7$ $38.5$ $39.05$
+DoRA $24.2$ $656.4$ $\mathbf{45.0}$ $\mathbf{47.2}$ $26.0$ $34.4$ $44.1$ $36.7$ $38.90$
+GaLore $23.7$ $437.7$ $42.2$ $45.3$ $28.0$ $39.0$ $43.1$ $41.2$ $39.80$
+LoSiA (-Pro) $21.9$ $\mathbf{290.4\ (244.8)}$ $44.7$ $46.7$ $\mathbf{28.4}$ $\mathbf{39.4}$ $\mathbf{45.0}$ $\mathbf{41.5}$ $\mathbf{40.95}$
+LLaMA 2-13B FFT-8Bit $77.1$ $640.7$ $61.2$ $55.7$ $35.7$ $43.2$ $53.6$ $56.2$ $50.93$
+LoRA $36.9$ $621.1$ $58.6$ $\mathbf{56.4}$ $34.1$ $44.8$ $52.6$ $53.7$ $50.03$
+PiSSA $36.9$ $622.3$ $53.4$ $55.2$ $34.5$ $44.8$ $52.0$ $48.8$ $48.11$
+LoSiA (-Pro) $36.9$ $\mathbf{548.5\ (453.4)}$ $\mathbf{59.0}$ $54.0$ $\mathbf{34.9}$ $\mathbf{48.2}$ $\mathbf{53.1}$ $\mathbf{55.7}$ $\mathbf{50.82}$
 Implementation Details
@@
 The training batch size is set to $4$, the warm-up ratio is set to $0.1$ and the model is trained by $3$ epochs.
-For training stability, the backbone model are in precision of BF16 and low-rank modules are upcasted to FP32.
+For training stability, the backbone models are in precision of BF16 and low-rank modules are upcasted to FP32.
 [^2] Trained with LLaMA-Factory Zheng et al.
 (2024).
-Upcasting to FP32 only costs an additional 0.6GB of memory and trains faster than BF16 in practice.
+Upcasting to FP32 only costs an additional 0.60.6GB of memory.
 All of the experiments are conducted on single NVIDIA A800 80GB GPU.
-Further details (including implementation details for common-sense reasoning tasks) can be found in Appendix A.3.
+Further details of the experimental setting (including implementation details on common-sense reasoning tasks) can be found in Appendix A.3.
 Table 2: Comparison of PEFT Methods on Commen-Sense Reasoning Tasks, using LLaMA-2 7B as the backbone model. Evaluations are PPL-based in lm-evaluation-harness and we report the ACC metric.
@@
 LoRA $19.46$ $10.0$ $50.28$ $79.71$ $59.86$ $73.88$ $79.33$ $55.00$ $56.86$ $\mathbf{88.07}$ $67.87$
-PiSSA $19.18$ $10.4$ $51.19$ $79.80$ $62.36$ $77.74$ $80.41$ $56.60$ $59.88$ $87.71$ $69.46$
+PiSSA $19.46$ $10.1$ $51.19$ $79.80$ $62.36$ $77.74$ $80.41$ $56.60$ $59.88$ $87.71$ $69.46$
 DoRA $20.42$ $25.6$ $51.71$ $79.34$ $59.86$ $\mathbf{79.24}$ $79.98$ $59.60$ $59.57$ $88.04$ $69.67$
@@
 4.2 Main Results
-Table 1 presents the performance of LoSiA compared to baseline methods across Gemma-2B, LLaMA2-7B, and LLaMA2-13B models.
+Table 1 presents the overall performance of LoSiA compared to baseline methods across Gemma-2B, LLaMA2-7B, and LLaMA2-13B models.
 For GSM8K, we report 0-shot Chain-of-Thought (CoT) and 5-shot accuracy to reveal the model’s reasoning capability and few-shot prompting performance.
@@
 The metrics are intended to measure the quality of generation and knowledge proficiency, respectively.
-Table 2 shows results on common-sense reasoning tasks, extracting the option with minimum perplexity and reporting ACC metric following lm-evaluation-harness.
-The test setup provides a robust measure of intrinsic knowledge acquisition.
+Table 2 shows the results on common-sense reasoning tasks, extracting the option with minimum perplexity, and reporting ACC metric following lm-evaluation-harness.
+The evaluation provides a robust measure of intrinsic knowledge acquisition.
 LoSiA effectively reserves knowledge
 LoSiA demonstrates superior knowledge retention, as evidenced by perplexity-based evaluations.
-It outperforms LoRA by 2.48% on commonsense reasoning tasks and maintains an average 1.93% improvement on MMLU (0-shot, PPL).
+It outperforms LoRA by 2.48% on common-sense reasoning tasks and maintains an average 1.93% improvement on MMLU (0-shot, PPL).
 Unlike low-rank methods, LoSiA’s sparse, high-rank fine-tuning approach enables localized knowledge retention while shifting likelihood toward correct answers.
@@
 Figure 5 compares the training overheads of various PEFT methods.
-Contrast to baselines such as DoRA which incur significant additional FLOPs, LoSiA shows superior efficiency in both training time and memory usage.
+Contrast to baselines such as DoRA that incur significant additional FLOPs, LoSiA shows superior efficiency in both training latency and memory usage.
 By eliminating extra matrix multiplication operations, LoSiA achieves faster training speeds.
-Its refined implement, LoSiA-Pro, further compresses activation storage by at least 22.8GB (w GC) and reduce the training time up to 34% (w/o GC) compared to LoRA by saving and computing on partial activations.
+Its refined implement, LoSiA-Pro, further compresses activation storage by at least 22.8GB (w/o GC) and raise training throughput to 1.38x (w GC) compared to LoRA by saving and computing on partial activations.
 A detailed training latency and GPU memory measurement is in Appendix A.4.
 4.3 Ablation Study
-This section assesses the functionality of sensitivity importance-aware localization, asynchronous mechanism, and re-warmups, alongside with robustness analysis of LoSiA.
+This section assesses the functionality of sensitivity importance-aware localization, asynchronous mechanism, re-warmings, and re-localizations, alongside with robustness analysis of LoSiA.
 We present comprehensive ablation studies in Table 3 and training dynamics in Figure 6.
 Additional robustness tests for rank factor selection are provided in the Appendix.
-Table 3: Ablation Study of LoSiA on GSM8K and MMLU, using LLaMA-2 7B as the backbone.
+Figure 6: Loss Curves of Baselines and LoSiA Variants, training on MetaMathQA and Alpaca-GPT4.
+Table 3: Ablation Study of LoSiA on GSM8K and MMLU benchmark, using LLaMA-2 7B as backbone.
 Model GSM8K MMLU Avg.
@@
 $43.96$ $44.32$ $44.14$
-Figure 6: Loss Curves of Baselines and LoSiA Variants, training on MetaMathQA and Alpaca-GPT4.
-Asynchronous mechanism yields more stable training
-Variant SL refers to using a synchronous layer-wise localization mechanism.
-However, it causes loss fluctuation, destabilizes later training, and degenerate the model performance by 1.36% in average, while asynchronous updates produce more stable loss curves.
+w/o Re-localization
+$42.76$ $43.81$ $43.29$
+Asynchronous re-localization yields more stable training
+When each layer fine-tunes with fixed core subnets as ReLO represents, it results in persistent under-fitting and marked drop in test accuracy, confirming that key parameters shift frequently during training, necessitating dynamic and periodic localization of the selected core sub-network to adapt in real time.
+Variant SL refers to using a synchronous layer-wise localization strategy.
+However, it causes loss fluctuation, destabilizes later training, and degenerates the model performance by 1.36% on average, while asynchronous updates produce more stable loss curves.
 Sensitivity-based importance versus gradient-based importance
-Variant GL uses absolute gradients as the importance score.
-On MMLU, its performance remains comparable but is biased toward Humanities tasks (see Table 11), while its accuracy on GSM8K drops by 1.66%.
-Sensitivity-based scores, which aggregate multi-sample information, are more effective to capture general patterns in linear layers compared to biased gradients.
-However, gradient-based LoSiA exhibits promising results.
-In practice, the storage of $\overline{I}(\cdot),\overline{U}(\cdot)$ (about 1GB memory occupation on LLaMA-2 7B) can be eliminated using gradient-based importance if needed.
+Variant GL adopts absolute gradients as the importance score.
+On MMLU, its performance remains comparable to LoSiA but is biased towards humanities tasks (see Table 12), while its accuracy on GSM8K drops by 1.66%.
+Sensitivity-based scores, which aggregate multi-sample information, are more effective in capturing general patterns in linear layers compared to biased gradients.
+However, the gradient-based variant exhibits promising results.
+In practice, the storage of $\overline{I}(\cdot),\overline{U}(\cdot)$ (about 1GB of memory occupation on LLaMA-2 7B) can be eliminated using gradient-based importance if needed.
 Further discussion is provided in Appendix A.2.2.
@@
 In permissible GPU memory constraints, fully training the output layer shows promising performance and is also recommended.
-Table 4: Robustness of Time Slot $T$ Across a Series of Data Scales. Trained by MetaMathQA and evaluated by GSM8K on LLaMA-2 7B.
+Table 4: Robustness of Time Slot $T$ Across a Series of Data Scales. Trained with MetaMathQA and evaluated by GSM8K on LLaMA-2 7B.
 Method @30K @50K @70K
@@
 Robustness across varying data scales and time slot lengths
-Table 4 assesses the performance of LoSiA across different training data scales.
-LoSiA consistently outperforms LoRA, demonstrating stability and robustness.
-Furthermore, the optimal time slot $T$ is positively correlated with the size of training set, while LoSiA shows transcendent performance within a reasonable range of $T$.
+Table 4 benchmarks LoSiA against LoRA as the training corpus grows.
+Across every scale, LoSiA consistently outperforms LoRA, demonstrating stability and robustness.
+Furthermore, the optimal time slot $T$ is positively correlated with the size of the training set, while LoSiA shows transcendent performance within a reasonable range of $T$.
 4.4 Analysis
 Selection Distribution
-We analyze the core subnet selection frequency of neurons in Figure 7.
-The frequently selected neurons remain similar under different rank factor $p$, while smaller $p$ produces more concentrated distribution patterns.
-This indicates that LoSiA effectively identifies and optimizes critical neurons with limited training budgets, while simultaneously adjusting marginal parameters to further enhance generalization capability.
+Figure 7 visualizes how often each neuron is chosen in core subnets during training.
+The frequently selected neurons remain similar under different rank factors $p$.
+The smaller $p$ merely sharpens the histogram: mass is pushed into the most important weights, so more radical compression does not discard salient parameters.
+LoSiA simultaneously adjusts marginal parameters to enhance generalization capability, yielding better generalization whenever the budget is tight.
+Figure 7: Selected Frequency Distributions of Neurons in Core Subnets. Sorted frequencies are ploted in black.
 Reduce Intruder Dimensions
@@
 Both LoRA and DoRA exhibit dimensional shifts due to their low-rank structures, whereas LoSiA demonstrates higher similarity and dimensional stability comparable to FFT.
-Figure 7: Selected Frequency Distributions of Neurons in Core Subnets. Sorted frequencies are ploted in black.
 Figure 8: Similarities of Top-500 Largest Singular Vectors between Pre- and Post-Fine-Tuning Weights.
 To evaluate LoSiA’s efficacy in continual learning, we perform sequential fine-tuning on Hellaswag, PiQA, BoolQ, SiQA, and Winogrande datasets on LLaMA-2 7B.
-We employs Average Performance (AP) Chaudhry et al.
-(2018), Forward Transfer (FWT) Lopez-Paz and Ranzato (2017), and Backward Transfer (BWT) Ke and Liu (2023) to assess overall performance, knowledge transfer ability from previous tasks to current task, and level of forgetting, respectively.
-Details of the metrics and experiments are provided in A.3.4.
+We employ Average Performance (AP) Chaudhry et al.
+(2018), Forward Transfer (FWT) Lopez-Paz and Ranzato (2017), and Backward Transfer (BWT) Ke and Liu (2023) metrics to assess overall performance, knowledge transfer ability from previous tasks to current task, and level of forgetting, respectively.
+Details of the experiments are provided in Appendix A.3.4.
+Suppose that the model learns sequentially on $N$ tasks.
+Let $P_{i,j}$ denote the accuracy on task $j$ after training on task $i$.
+Following Zhao et al.
+(2024b); Feng et al.
+(2025), we formulate the metrics (AP, FWT and BWT) as bellow:
+Average Performance: The metric reflects overall task performance after continued learning, which is, $AP=\frac{1}{N}\sum_{i=1}^{N}P_{N,i}$
+Forward Transfer: The metric measures the transferability of learned knowledge from previous tasks to a new task.
+$\text{FWT}=\frac{1}{N}\sum_{i=1}^{N}(P_{i,i}-P_{0,i})$, where $P_{0,i}$ is the performance of individually training task $i$.
+Backward Transfer: The metric evaluates the impact of learning later tasks on the model’s performance on an earlier task, that is, $\text{BWT}=\frac{1}{N-1}\sum_{i=1}^{N-1}(P_{N,i}-P_{i,i})$.
 Table 5: Results of Continue Learning with Sequential PEFTs on Five Commen-Sense Reasoning Tasks.
@@
 Extensive experiments show that LoSiA outperforms baselines on domain-specific and common-sense reasoning tasks while reducing forgetting.
-We hope that our work will inspire future research to further explore the intrinsic substructures in supervised fine-tuning.
+We hope that our work will inspire future research to further explore intrinsic substructures in supervised fine-tuning.
 6 Limitation
@@
 Moreover, to accommodate diverse datasets and practical deployment conditions, automated time slot selection mechanisms warrant further investigation.
+Acknowledgements
+This work is supported by the National Natural Science Foundation of China (No.
+52539001).
+It’s partially supported by CHN Energy Dadu River Big Data Services Co., Ltd.
+It’s also supported by Student Research Training (SRT) project of Tsinghua University.
+We also thank anonymous reviewers for their valuable feedback.
 Appendix A Appendix
@@
 A.1.1 Proof for Formula 1
-On a batch $\mathcal{B}$ composed of $M$ samples, the MSE Loss between full fine-tuning (produces model $f$) and training with parameter set $P$ (produces model $f^{\prime}$) is given by:
+On a batch $\mathcal{B}$ composed of $M$ samples, the MSE loss between full fine-tuning (which produces model $f$) and training on parameter set $P$ (which produces model $f^{\prime}$) is given by:
 $\mathcal{L}_{MSE}$ $=\frac{\|y-y^{\prime}\|^{2}_{F}}{M}=\frac{\|Wx-W^{\prime}x\|^{2}_{F}}{M}$ (10)
@@
 SGD
-In SGD optimizer, suppose learning rate is $\eta$, the difference in fine-tuned parameter is:
-$W-W^{\prime}=-\eta\mathbb{1}_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B})$ (12)
+In SGD optimizer, supposing the learning rate is $\eta$, the difference in fine-tuned parameter is:
+$W-W^{\prime}=-\eta 1_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B})$ (12)
 It derives an upper bound for the MSE Loss:
-$\mathcal{L}_{MSE}$ $\leq\eta^{2}\frac{\|\mathbb{1}_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B})\|^{2}_{F}\|x\|^{2}_{F}}{M}$ (13)
-The result suggests that maximizing the sum of $\nabla_{W_{0}}\mathcal{L}(\mathcal{B})_{ij}$ where $(i,j)\in P$ ideally tightens the approximate error of training on subset.
+$\mathcal{L}_{MSE}$ $\leq\eta^{2}\frac{\|1_{(i,j)\not\in P}\cdot\nabla_{W_{0}}\mathcal{L}(\mathcal{B})\|^{2}_{F}\|x\|^{2}_{F}}{M}$ (13)
+The result suggests that maximizing the sum of $\nabla_{W_{0}}\mathcal{L}(\mathcal{B})_{ij}$ where $(i,j)\in P$ ideally tightens the approximate error of training on parameter subset.
 AdamW
@@
 $\tilde{G_{t}}$= $\frac{M_{t}}{\sqrt{V_{t}+\epsilon}}$ (17)
-Similarly, since $W-W^{\prime}=-\eta\mathbb{1}_{(i,j)\not\in P}\cdot\tilde{G_{t}}$, we analysis relationship between $\tilde{G_{t}}$ and $G_{t}$:
+Similarly, since $W-W^{\prime}=-\eta 1_{(i,j)\not\in P}\cdot\tilde{G_{t}}$, we analyze the relationship between $\tilde{G_{t}}$ and the original gradient $G_{t}$ by element:
 $\begin{split}\frac{\partial(\tilde{G_{t}})^{2}}{\partial G_{t}}=&2M_{t}[\frac{(1-\beta_{1})V_{t}}{V_{t}^{2}}\\ &-\frac{(1-\beta_{2})G_{t}M_{t}}{V_{t}^{2}}]\end{split}$ (18)
 Suppose $M_{t}>0$, when $G_{t}<\frac{(1-\beta_{1})V_{t}}{(1-\beta_{2})M_{t}}$, $\frac{\partial(\tilde{G_{t}})^{2}}{\partial G_{t}}>0$.
-In practice, the typical settings are $\beta_{1}=0.9,\beta_{2}=0.999$.
+In practice, typical settings are $\beta_{1}=0.9,\beta_{2}=0.999$.
 Therefore, when $G_{t}<10^{2}\frac{V_{t}}{M_{t}}$, $\tilde{G}_{t}$ increases with $G_{t}$, effectively covering a broad range of non-stationary optimization scenarios.
@@
 (2016).
-However, for real-time importance calculation during training, approximations is necessary and is derived bellow.
+However, to establish real-time importance calculation during training, approximations are necessary and are derived below.
 Element-wise importance score $I(\cdot)$ is formulated as:
-$\begin{split}I(W_{k})=&|\Delta\mathcal{L}(\mathcal{D})|=|\mathcal{L}(\mathcal{D})-\mathcal{L}_{W_{k}=0}(\mathcal{D})|\\ =&|\frac{\partial\mathcal{L}^{T}(\mathcal{D})}{\partial{W_{k}}}W_{k}-\frac{1}{2}W_{k}^{T}H_{kk}W_{k}\\ &+o(W_{k}^{2})|\end{split}$ (19)
-where $H$ denotes the Hessian matrix. It is computational intensive for Hessian calculation, we therefore use fisher information matrix $F$ for diagonal elements instead:
+$\begin{split}I(W_{k})=&|\Delta\mathcal{L}(\mathcal{D})|=|\mathcal{L}(\mathcal{D})-\mathcal{L}_{W_{k}=0}(\mathcal{D})|\\ =&|\frac{\partial\mathcal{L}^{T}(\mathcal{D})}{\partial{W_{k}}}W_{k}-\frac{1}{2}W_{k}H_{kk}W_{k}\\ &+o(W_{k}^{2})|\end{split}$ (19)
+where $H$ denotes the Hessian matrix, which is computationally intensive. Therefore, the fisher information matrix $F$ is used instead to obtain diagonal elements of the Hessian matrix:
 $\begin{split}F_{kk}=&-H_{kk}=-\mathbb{E}_{p(\theta|\mathcal{D})}[\frac{\partial^{2}\mathcal{L}(\theta,D)}{\partial^{2}\theta_{k}}|_{\theta=\theta^{*}}]\\ \approx&-\mathbb{E}_{(x,y)\sim\mathcal{D}}[(\frac{\partial\mathcal{L}(\theta,x,y)}{\partial\theta_{k}}|_{\theta=\theta^{*}})^{2}]\end{split}$ (20)
-Approximating the expectation by dataset $\mathcal{D}$ based on Monte Carlo Method, it derives:
+Approximating mathematical expectations on dataset $\mathcal{D}$ using the Monte Carlo method derives:
 $\begin{split}I(W_{k})=&|\frac{\partial\mathcal{L}(\mathcal{D})}{\partial{W_{k}}}W_{k}+o(W_{k}^{2})\\ &-\sum_{(x,y)\in\mathcal{D}}\frac{1}{2|\mathcal{D}|}(\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}})^{2}W_{k}^{2}|\end{split}$ (21)
 During training, the dataset $\mathcal{D}$ is processed in batches $\mathcal{B}_{i}$, and the batch gradient is calculated as $\nabla_{W}\mathcal{L}(\mathcal{B}_{i})=\frac{1}{M}\sum_{j=1}^{M}\nabla_{W}\mathcal{L}(\mathcal{B}_{ij})$.
-To avoid calculating the gradients separately for each sample in the batch, we approximate $\sum_{(x,y)\in\mathcal{B}_{i}}\frac{1}{M}(\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}})^{2}$ to the term $(\frac{\sum_{(x,y)\in\mathcal{B}_{i}}\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}}}{M})^{2}$.
+To avoid calculating the gradients separately for each sample in the batch, we approximate $\sum_{(x,y)\in\mathcal{B}_{i}}\frac{1}{M}(\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}})^{2}$ to the term of $(\frac{\sum_{(x,y)\in\mathcal{B}_{i}}\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}}}{M})^{2}$.
 To analyze errors, assume $g=\frac{\partial\mathcal{L}(x,y)}{\partial{W_{k}}}\sim G$, we have:
@@
 Task
-Given an arbitrary matrix $A^{n\times m}$, select $\tilde{n}\ (\tilde{n}\leq n)$ rows $X_{S}$ and $\tilde{m}\ (\tilde{m}\leq m)$ columns $Y_{S}$, to maximize the sum $\sum_{i\in X_{S}}\sum_{j\in Y_{S}}A_{ij}$.
+Given an arbitrary non negative matrix $A^{n\times m}$ and cardinal budget requirements $\tilde{n},\tilde{m}$.
+Select $\tilde{n}$ rows $X_{S}$ and $\tilde{m}$ columns $Y_{S}$ to maximize the sum $\sum_{i\in X_{S}}\sum_{j\in Y_{S}}A_{ij}$.
 Lemma
@@
 Proof
-Construct a special form of $A^{n\times n}$ as the adjacent matrix of graph $G$ with larger values on diagonal maximum, that is:
+Construct a special form of $A^{n\times n}$ as the adjacent matrix of graph $G$ with larger values on the diagonal maximum, that is:
 $A_{uv}=\begin{cases}1&\text{if }(u,v)\in E,\\ n^{2}+1&\text{if }u=v\\ 0&\text{otherwise}\end{cases}$
-Then, MCP can be reduced to the task in polynomial time following the algorithm:
+Then, the MCP problem can be reduced to the task in polynomial time following the algorithm below:
   • Enumerate $k$ in descending order $n,n-1\ldots 1$
@@
   • If the optimal solution equals to $(n^{2}+k)k$, then there exist a clique $C=X_{S}$ of size $k$, terminate.
-Therefore, a NP-Complete problem can be reduced to the task in polynomial time, which yields the conclusion that the task is NP-Hard.
+Therefore, an NP-Complete problem can be reduced to the task in polynomial time, which yields the conclusion that the task is NP-Hard.
 A.2 Further Observations
@@
 To investigate the universality of the sparse sub-network structure for large gradients, we analyze gradient magnitude distributions across different layers, as shown in Figure 9.
-Both Gradients of the Self-Attention and MLP modules exhibits the consistent structure of core subnets.
+Both Gradients of the Self-Attention and MLP modules exhibit the consistent structure of core subnets.
+We also quantify this observation on different layers and each submodule in Table 6.
+Core-subnet localization markedly outperforms random selection and closely approaches the ideal Top-K baseline, corroborating the validity of sparse-gradient patterns.
+Note, however, that the ideal Top-K set is irregular and therefore entails nontrivial runtime overhead.
+Table 6: Sum of Absolute Gradient Values ($\times 10^{3}$) for Selection Patterns. "Subnet" stands for the core subnet localization algorithm, while Top-K is an ideal selection.
+Layer Submodule Total Random Subnet Top-K (Ideal)
+5 q_proj $16.38$ $1.02$ $3.82$ $6.72$
+k_proj $14.02$ $0.88$ $3.18$ $5.70$
+v_proj $83.97$ $5.25$ $18.69$ $30.59$
+o_proj $93.18$ $5.82$ $19.20$ $29.70$
+up_proj $133.12$ $8.32$ $20.48$ $33.02$
+down_proj $144.38$ $9.02$ $21.89$ $37.12$
+gate_proj $101.89$ $6.37$ $16.51$ $27.26$
+15 q_proj $15.49$ $0.97$ $4.83$ $6.62$
+k_proj $12.29$ $0.77$ $3.22$ $5.25$
+v_proj $48.64$ $3.04$ $12.93$ $16.90$
+o_proj $48.38$ $3.02$ $12.29$ $14.98$
+up_proj $65.02$ $4.06$ $15.30$ $20.86$
+down_proj $70.14$ $4.38$ $15.42$ $21.63$
+gate_proj $54.53$ $3.41$ $13.63$ $19.07$
+25 q_proj $6.94$ $0.43$ $2.82$ $4.13$
+k_proj $6.02$ $0.38$ $1.90$ $3.54$
+v_proj $12.93$ $0.81$ $4.10$ $5.98$
+o_proj $13.38$ $0.84$ $4.61$ $5.86$
+up_proj $32.51$ $2.03$ $8.13$ $10.82$
+down_proj $36.10$ $2.26$ $8.64$ $11.20$
+gate_proj $25.09$ $1.57$ $6.50$ $8.70$
 (a) Decoder Layer $15$
@@
 The datasets of common-sense reasoning tasks are presented in Table 9, while corresponding hyperparameters detailed in Table 10.
-The GPU memory usage remain aligned.
+The GPU memory usage remains aligned.
 For each PEFT baselines, searches in learning rate are performed.
@@
 A.3.3 Rank Factor Robustness
-To evaluate the impact of the rank factor $p$, which determines the scale of the core subnets, we conduct an ablation study on MetaMathQA.
+To evaluate the impact of the rank factor $p$, which determines the scale of selected core subnets, we conduct an ablation study on MetaMathQA as Table 11 shows.
+Performance grows steadily with the number of training parameters on both Gemma 2B and LLaMA-2 7B.
 The results demonstrate LoSiA’s robustness across various subnet scales.
-Note that $p=1/16$ may be relatively small for effective subnet fine-tuning, while increasing computational budget boosts the performance.
-Table 10: Rank Factor Robustness on GSM8K
+Note that $p=1/16$ may be relatively small for effective subnet fine-tuning, while increasing the computational budget boosts the performance.
+Table 11: Rank Factor Robustness of LoSiA on GSM8K
 Model $\mathbf{1/16}$ $\mathbf{1/8}$ $\mathbf{1/4}$ $\mathbf{1/2}$
-Gemma-2B $37.53$ $42.84$ $45.03$ $45.64$
+Gemma 2B $37.53$ $42.84$ $45.03$ $45.64$
 LLaMA-2 7B $40.64$ $44.66$ $46.02$ $48.45$
@@
 A.3.4 Continue Learning
-To examine whether reduction of intruder dimensions in LoSiA mitigates forgetting in continue learning, we sequentially adapt LLaMA-2 7B through five common-sense reasoning tasks by the order HellaSwag, PIQA, BoolQ, SIQA and Winogrande.
+To examine whether reduction of intruder dimensions in LoSiA mitigates forgetting in continue learning, we sequentially adapt LLaMA-2 7B through five common-sense reasoning tasks: HellaSwag, PIQA, BoolQ, SIQA and Winogrande.
 Learning rate for LoRA is 1e41e-4 and for LoSiA is 5e55e-5.
@@
 LoRA modules are merged into the backbone before subsequent task adaptation.
-Suppose the model learn sequentially on $N$ tasks.
-Let $P_{i,j}$ denote the ACC on task $j$ after training on task $i$.
-Following Zhao et al.
-(2024b), we formulate the metrics (AP, FWT and BWT) as bellow:
-Average Performance: $AP=\frac{1}{N}\sum_{i=1}^{N}P_{N,i}$
-Forward Transfer: The metric measures the transferability of learned knowledge from previous tasks to a new task.
-$\text{FWT}=\frac{1}{N}\sum_{i=1}^{N}(P_{i,i}-P_{0,i})$, where $P_{0,i}$ is the performance of individually training task $i$.
-Backward Transfer: The metric evaluates the impact of learning later tasks on the model’s performance on an earlier task, that is $\text{BWT}=\frac{1}{N-1}\sum_{i=1}^{N-1}(P_{N,i}-P_{i,i})$.
 Table 13 shows the detailed result during sequential adaptation.
@@
 Table 13: Details of Performances on Continue Learning Five Common-Sense Reasoning Tasks. The column stands for training order, while the label "ST" indicates the result in single-tasking training.
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
+Method Task (#1) HellaS (#2) PIQA (#3) BoolQ (#4) SIQA (#5) WinoG ST
+Seq-LoRA HellaSwag $59.86$ $55.64$ $59.10$ $57.86$ $54.36$ $59.86$
+PIQA $76.01$ $80.52$ $77.86$ $78.73$ $77.64$ $79.33$
+BoolQ $77.80$ $73.27$ $86.30$ $80.12$ $75.93$ $88.07$
+SIQA $45.80$ $47.80$ $45.85$ $59.52$ $46.11$ $56.86$
+Winogrande $64.25$ $68.35$ $68.82$ $69.93$ $79.08$ $73.88$
+Seq-LoSiA HellaSwag $63.72$ $61.89$ $61.11$ $60.37$ $\mathbf{56.43}$ $63.72$
+PIQA $78.29$ $79.49$ $79.82$ $79.38$ $\mathbf{77.75}$ $81.50$
+BoolQ $77.52$ $70.76$ $83.24$ $82.54$ $\mathbf{81.99}$ $84.13$
+SIQA $47.80$ $48.26$ $48.26$ $59.93$ $\mathbf{56.04}$ $61.05$
+Winogrande $68.51$ $67.88$ $68.51$ $71.82$ $\mathbf{80.19}$ $77.19$
 A.4 Resources Measurement
@@
 Table 14 shows GPU memory consumption details of LoRA, GaLore and LoSiA.
+For optimizers like AdamW, LoSiA reduces the gradient dimension of the output layer to a fraction $p_{o}$, while GaLore performs full fine-tuning on the output layer of shape $d\times V$.
+Both GaLore and LoSiA utilize per-layer weight update techniques for gradient computation.
+The update is therefore computed upon gradient acquisition and then promptly discarded.
+In terms of auxiliary parameters, GaLore requires storing down- and up-projection matrices.
+Since $R$ is typically high-rank, GaLore’s auxiliary parameters can be significantly larger than those of other methods, which may induce additional GPU memory consumptions.
+For LoSiA, auxiliary parameters are used to compute the importance scores ($\overline{U}(\cdot)$ and $\overline{I}(\cdot)$).
+If gradient-based importance scoring is adopted, this component can be completely eliminated.
+Regarding total memory consumption, increasing the rank in LoRA and GaLore incurs substantial overhead.
+However, in LoSiA, only the term $2(LKd^{2}p^{2}b+Vdp_{o}b)$ scales with rank factor $p$.
+When using LoSiA-Pro without Gradient Checkpointing, only the activations corresponding to the input neurons need to be stored, making it the sole PEFT approach capable of mitigating this class of memory bottlenecks.
 Table 14: Comparison of Memory Consumptions.
-Cells in green highlight the components that may notably lower than other methods, while in red highlight the components that may cause relatively large memory consumption.
+Cells in green highlight the components that may notably lower than other methods, while in red highlight the components that may cause relatively large memory consumption in high-rank.
 LoRA GaLore LoSiA
@@
 $2Kd^{2}b$
-For optimizers like AdamW, LoSiA reduces the gradient dimension of the output layer to a fraction $p_{o}$, while GaLore performs full fine-tuning on the output layer of shape $d\times V$.
-Both GaLore and LoSiA utilize per-layer weight update techniques for gradient computation.
-In contrast, basic LoRA implementations require collecting all gradients after backward propagation.
-The reduced memory overhead of gradient storage allows GaLore and LoSiA to fully train the $\mathtt{lm\_head}$ in LLaMA-2 models.
-In terms of auxiliary parameters, GaLore requires storing down- and up-projection matrices.
-Since $R$ is typically high-rank, GaLore’s auxiliary parameters can be significantly larger than those of other methods.
-For LoSiA, auxiliary parameters are used to compute the importance scores ($\overline{U}(\cdot)$ and $\overline{I}(\cdot)$).
-If gradient-based importance scoring is adopted, this component can be completely eliminated.
-Regarding total memory consumption, increasing the rank in LoRA and GaLore incurs substantial overhead.
-However, in LoSiA, only the term $2(LKd^{2}p^{2}b+Vdp_{o}b)$ scales with rank factor $p$, resulting in a more efficient memory footprint.
-Table 14: Details of Trainable Parameters for LoRA and LoSiA under Different Hyperparameter Configurations on LLaMA-2 7B.
-LoRA
-$\mathbf{Rank\ r}$ $16$ $64$ $256$ $1024$
-$\mathbf{\#Trainable}$ 40.040.0M 160.0160.0M 640.0640.0M 2560.02560.0M
-$\mathbf{Mem(GB)}$ $22.33$ $23.72$ $28.42$ $63.90$
+Table 15: Details of Trainable Parameters for LoSiA under Different Hyperparameter Configurations on LLaMA-2 7B.
 LoSiA
@@
 A.4.2 Latency Measurement
-We measure the training latency ($\mu s$ / token) fine-tuning with different PEFT methods on LLaMA-2 7B, and the results are shown in Table 15.
-The experiments are conducted with $\text{cutoff\_len}=2048$ and $\text{batch \_size}=4$.
 Table 16: Comparison of Training Latency on LLaMA-2 7B. Latencies are reported in measuerments of $\mu s$ per token, training with Flash-Attention 2 Dao (2023).
 Forward Backward Other Total
-w Gradient Check-Pointing
+w Gradient Checkpointing
 LoRAr=64 74.0 264.0 0 338.0
@@
 […]
-w/o Gradient Check-Pointing
+w/o Gradient Checkpointing
 LoRAr=64 Out of Memory
@@
 […]
+We measure the training latency ($\mu s$ / token) fine-tuning with different PEFT methods on LLaMA-2 7B, and the results are shown in Table 16.
+The experiments are conducted with $\text{cutoff\_len}=2048$ and $\text{batch \_size}=4$.
 While demonstrating superior performance among existing baselines, LoSiA reduces training latency by 14.1% compared to LoRA, 55.8% compared to DoRA.
@@
 Specifically, during backward propagation with Gradient Checkpointing, the production of low-rank matrices introduces significant overhead for activation recomputation and gradient calculation.
-Note that LoRA can avoid gradient calculations on backbone weights, but this requires specialized implementations and introduces a large coefficient of computational complexity.
+Note that LoRA can avoid gradient calculations on backbone weights, but this requires specialized implementations and still introduces a large coefficient of computational complexity.
 For LoSiA-Pro, the computational complexity remains the same as LoSiA during the forward pass, but it only requires storing a proportion $p$ of the input activations of the linear layers.
@@
 This results in highly efficient training and lower GPU memory consumption.
+A.4.3 Algorithm
+The core of LoSiA is summarized in Algorithm 2.
+The model is partitioned into weight groups (in this paper, simply the decoder layers); each group is assigned its own LoSiA optimizer that receives the total number of groups ($L$) and related meta-information, and LoSiA automatically performs the weight updates during the backward pass.
+Algorithm 2 Pseudo Code of LoSiA
+1:Weight matrix $W\in\mathbb{R}^{n\times m}$ lying in decoder layer (weight groups) $l$; Total decoder layer (weight groups) $L$; EMA ratio $\beta_{1},\beta_{2}$; Adam decay rates $\beta^{\prime}_{1},\beta^{\prime}_{2}$; Rank factor $p$; Time slot $T$;
+2:
+3:Initialize scales of the core subnet $n_{p}\leftarrow\lfloor np\rfloor,m_{p}\leftarrow\lfloor mp\rfloor$
+4:Initialize first- and second-order momentum $M_{0}\leftarrow 0_{n_{p}\times m_{p}}$, $V_{0}\leftarrow 0_{n_{p}\times m_{p}}$
+5:Initialize selected neurons randomly $\rho\leftarrow\text{random}([1\ldots n],n_{p})$, $\gamma\leftarrow\text{random}([1\ldots m],m_{p})$
+6:Initialize training step $t\leftarrow 1$
+7:
+8:repeat
+9:  $t^{\prime}\leftarrow(t-1)\mod(TL)$
+10:  if $\lfloor\frac{t^{\prime}}{T}\rfloor=l-1$ then $\triangleright$ {sensitivity-based parameter importance estimation}
+11:   $I\leftarrow W\cdot\nabla W$
+12:   $I\leftarrow|I-\frac{1}{2}I^{2}|$ $\triangleright$ {calculation of importance score by Eq.3}
+13:
+14:   if $t$ is the first step of time slot $T$ then
+15:     $\overline{I}_{t-1}\leftarrow 0_{n\times m}$, $\overline{U}_{t-1}\leftarrow 0_{n\times m}$
+16:   end if
+[…]
+18:   $\overline{I}_{t}\leftarrow\beta_{1}\overline{I}_{t-1}+(1-\beta_{1})\overline{I}_{t}$
+19:   $\overline{U}_{t}\leftarrow\beta_{2}\overline{U}_{t-1}+(1-\beta_{2})\overline{U}_{t}$
+20:  end if
+21:  WEIGHTS OPTIMIZATION BY Adam
+22:   ($\nabla W_{\rho,\gamma}\leftarrow\tilde{L}_{S}\tilde{R}_{S}$) $\triangleright$ {partial activation $\tilde{L}_{S}$ manually saved during forward in LoSiA-Pro}
+23:   $G\leftarrow\nabla W_{\rho,\gamma}$ $\triangleright$ {obtain subnet gradient by indices selection}
+24:   $M_{t}\leftarrow\beta^{\prime}_{1}M_{t-1}+(1-\beta^{\prime}_{1})\cdot G$
+25:   $V_{t}\leftarrow\beta^{\prime}_{2}V_{t-1}+(1-\beta^{\prime}_{2})\cdot G^{2}$
+26:   $M_{t}\leftarrow M_{t}/(1-\beta^{\prime}_{1})$, $V_{t}\leftarrow V_{t}/(1-\beta^{\prime}_{2})$
+27:   $N_{t}\leftarrow M_{t}/(\sqrt{V_{t}}+\epsilon)$
+28:
+29:   $\eta\leftarrow\overline{lr}(t)$ $\triangleright$ {calculate learning rate using rewarming scheduler in Section 3.3}
+30:   $W_{t}\leftarrow W_{t-1}-\eta N_{t}$
+31:
+32:
+33:  if $t^{\prime}\mod T=0$ and $\frac{t^{\prime}}{T}=l$ then
+34:   $s_{W}\leftarrow\overline{I}_{t-1}\cdot\overline{U}_{t-1}$
+35:   for localization algorithm $\mathcal{A}_{i}$ do
+36:     $\rho_{i},\gamma_{i}\leftarrow\mathcal{A}_{i}(s_{W},p)$
+37:   end for
+38:   $k=\arg\max_{i}S(s_{W_{\rho_{i},\gamma_{i}}})$ $\triangleright$ {choose the core subnet with the highest importance}
+39:   $\rho,\gamma\leftarrow\rho_{k},\gamma_{k}$
+40:   $\text{delete}\ \overline{I},\overline{U}$ $\triangleright$ {delete useless tensors after importance calculation terminate}
+41:   $M_{t}\leftarrow 0_{n_{p}\times m_{p}},V_{t}\leftarrow 0_{n_{p}\times m_{p}}$
+42:  end if
+43:
+44:  $t\leftarrow t+1$
+45:until training finishes
+46:return $W$